### v1

# Lifting Generators in Connected Lie Groups

Abstract
Given an epimorphism between topological groups $f:G\to H$, when can a generating set of $H$ be lifted to a generating set of $G$?
We show that for connected Lie groups the problem is fundamentally abelian: generators can be lifted if and only if they can be lifted in the induced map between the abelianisations (assuming the number of generators is at least the minimal number of generators of $G$). As a consequence, we deduce that connected perfect Lie groups satisfy the Gaschütz lemma: generating sets of quotients can always be lifted. If the Lie group is not perfect, this may fail. The extent to which a group fails to satisfy the Gaschütz lemma is measured by its Gaschütz rank, which we bound for all connected Lie groups, and compute exactly in most cases. Additionally, we compute the maximal size of an irredundant generating set of connected abelian Lie groups, and discuss connections between such generation problems with the Wiegold conjecture.
1 Introduction
Many problems and results in group theory are concerned with generating sets of groups, and with lifting generators through quotients. A classical example is the Gaschütz lemma, which states the following: given an epimorphism between finite groups $f:G\to H$, every generating set $h_{1},...,h_{n}$ of $H$ can be lifted via $f$ to a generating set $g_{1},...,g_{n}$ of $G$, provided that $G$ can be generated by $n$ elements. This result has proved to be very useful in the study of finite groups and the automorphism group of the free group $\mathrm{Aut}(F_{n})$. Furthermore, its extension to profinite groups serves as a key tool in the study of presentations of groups, subgroup growth, and the Wiegold conjecture (which we further discuss below), see [Lub01, LS12, GL09, BCL22].
Generator lifting problems may be studied for general topological [^1] Unless otherwise stated, we assume in this section that all groups are Hausdorff. groups. In such case, generation is understood in the topological sense: a subset $S$ of a group $G$ is said to generate $G$ if the group abstractly generated by $S$ (which is denoted by $\left\langle S\right\rangle$) is dense in $G$. Following [CG18], we define
Definition 1.1.
A topological group $G$ is said to be Gaschütz if, whenever $f:G\to H$ is a continuous open [^2] The restriction to open maps is important when dealing with topological groups which admit weaker group topologies. epimorphism and $h_{1},...,h_{n}\in H$ generate $H$ with $n\geqslant\mathrm{d}(G)$, there exist lifts $g_{1},...,g_{n}\in G$ (i.e., elements satisfying $f(g_{i})=h_{i}$) that generate $G$.
Here and thereafter, $\mathrm{d}(G)$ denotes the minimal size of a generating set of $G$.
In the terminology of Definition 1.1, the Gaschütz lemma states that finite groups are Gaschütz [Gas55]. In [CG18], it is shown that all first-countable [^3] It is in fact enough to assume the connected component of the identity is first-countable. compact groups are Gaschütz as well. Our first result welcomes another class of groups to this family:
Theorem 1.2.
Every perfect connected Lie group is Gaschütz.
This includes, of course, all connected semisimple Lie groups. The assumption that $G$ is perfect (i.e., that the derived subgroup $G^{\prime}$ of $G$ is dense) is essential. In fact, any connected abelian Lie group which is not compact and not isomorphic to $\mathbb{R}$ is not Gaschütz (see 3.10). We demonstrate that this is the only source of difficulty, in the sense that the challenge of lifting generators is ultimately reduced to the abelian case. This is the content of our main theorem, from which Theorem 1.2 is an immediate consequence.
Theorem 1.3.
Let $G$ be a connected Lie group, $f:G\to H$ an open epimorphism. Let $f^{\mathrm{ab}}:G/\overline{G^{\prime}}\to H/f\left(\overline{G^{\prime}}\right)$ be the map [^4] Observe that $f\left(\overline{G^{\prime}}\right)$ might not be closed in $H$, so $H/f\left(\overline{G^{\prime}}\right)$ might not be Hausdorff. induced by $f$. If $h_{1},\dots,h_{n}\in H$ generate $H$ with $n\geqslant\mathrm{d}(G)$, then they admit generating lifts to $G$ via $f$ if and only if $h_{1}f\left(\overline{G^{\prime}}\right),\dots,h_{n}f\left(\overline{G^{\prime}}\right)$ admit generating lifts to $G/\overline{G^{\prime}}$ via $f^{\mathrm{ab}}$.
We are thus left with understanding the abelian case. As mentioned above, even $\mathbb{R}\times\mathbb{R}$ is not Gaschütz: it is generated by $3$ elements but the quotient $\mathbb{R}\times\mathbb{R}/\mathbb{Z}$ admits $3$ generators that cannot be lifted. However, every $4$ generators can be lifted. We therefore introduce the following notion.
Definition 1.4.
Let $f:G\to H$ be a continuous epimorphism between topological groups and let $n\in\mathbb{N}$. We say that $f$ is $n$-Gaschütz if for every $n$ elements $h_{1},...,h_{n}\in H$ that generate $H$ there are lifts $g_{1},...,g_{n}\in G$ that generate $G$.
The Gaschütz rank of $G$, denoted $\text{\lightning}(G)$, is the minimal $m\in\mathbb{N}\cup\left\{\infty\right\}$ such that every open epimorphism $f:G\to H$ is $n$-Gaschütz for every $n\geqslant m$.
Interestingly, we do not know if every $n$-Gaschütz map is necessarily $(n+1)$-Gaschütz. This seems to be an interesting question.
Obviously, $\mathrm{d}(G)\leqslant\text{\lightning}(G)$ holds for all groups. By definition, a group is Gaschütz precisely when the equality $\mathrm{d}(G)=\text{\lightning}(G)$ holds. The abelian case is settled by the following.
Theorem 1.5.
If $A$ is a connected abelian Lie group with maximal torus $T$, then
$\text{\lightning}(A)=\begin{cases}1&A\text{ is compact,}\\ 2\dim A-\dim T&A\text{ is non-compact.}\end{cases}$
We are thus able to estimate the Gaschütz rank of arbitrary connected Lie groups.
Corollary 1.6.
Let $G$ be a connected Lie group. Then
$\text{\lightning}(G)\leq\max\{\mathrm{d}(G),2\dim(G/\overline{G^{\prime}})-\dim T\}$
where $T$ is the maximal torus of the abelianisation $G/\overline{G^{\prime}}$. If $G/\overline{G^{\prime}}$ is non-compact, then equality holds.
The value $\mathrm{d}(G)$ of a connected Lie group $G$ was computed explicitly in [AN24]; see Section 6 for details. Thus, Corollary 1.6 provides an explicit bound for the Gaschütz rank. It moreover provides a precise value for the Gaschütz rank for most connected Lie groups. Indeed, the only case which remains undetermined is when $G$ is not compact, $G/\overline{G^{\prime}}$ is compact, and in addition $\dim G/\overline{G^{\prime}}>\mathrm{d}(G)$. The nature of this unique situation is quite subtle. By Corollary 1.6, $\text{\lightning}(G)$ is at most $\max\{\mathrm{d}(G),2\dim(G/\overline{G^{\prime}})-\dim T\}$ and at least $d(G)$. Both ends of the spectrum are possible; see 8.
Corollary 1.6 in particular states that the Gaschütz rank of any connected Lie group is finite. We note that this is not true for general groups. Indeed, there are connected locally compact groups (e.g., $G=\mathbb{R}\times(\mathbb{R}/\mathbb{Z})^{\aleph_{0}}$) with $\mathrm{d}(G)<\infty$ but $\text{\lightning}(G)=\infty$ However, we do not know the answer to the following question:
Question 1.7.
Does there exist a finitely generated discrete group with infinite Gaschütz rank?
The latter is equivalent to asking whether the free group $F_{n}$ on $n$ generators has a finite Gaschütz rank for every $n$.
The Wiegold Conjecture.
The Gaschütz rank of an epimorphism from $F_{n}$ onto a finite group is related to the Wiegold conjecture, which we now recall. For a discrete group $G$, denote by $\mathcal{G}_{n}(G)$ the space of epimorphisms $F_{n}\to G$, or equivalently the space of generating $n$-tuples of $G$. We have a natural action of $\mathrm{Aut}(F_{n})$ on $\mathcal{G}_{n}(G)$ by precomposition. The Wiegold conjecture states that the action of $\mathrm{Aut}(F_{3})$ on $\mathcal{G}_{3}(G)$ is transitive for every finite simple group $G$. A stronger conjecture, due to Pak [Pak01], states that the action of $\mathrm{Aut}(F_{n})$ on $\mathcal{G}_{n}(G)$ is transitive whenever $G$ is a finite group that can be generated by $n-1$ elements. See [Lub11] for more details.
It is easy to see the following:
Lemma 1.8.
Let $G$ be a discrete group. The action of $\mathrm{Aut}(F_{n})$ on $\mathcal{G}_{n}(G)$ is transitive if and only if every epimorphism $F_{n}\to G$ is $n$-Gaschütz.
Therefore, while it is known that $F_{3}$ is not Gaschütz, the question whether every epimorphism from $F_{3}$ onto a finite simple group is $3$-Gaschütz is a major open problem.
In light of this, one readily sees that the analogous question for profinite groups has a positive solution. That is, the action of $\mathrm{Aut}(\hat{F}_{n})$ on the space of epimorphisms $\hat{F}_{n}\to G$ is always transitive, since the free profinite group $\hat{F}_{n}$ is Gaschütz. This is a key idea in [Lub01, BCL22] and in [GL09, §2.3].
When $G$ is a finite simple group and $n=2$, the action of $\mathrm{Aut}(F_{2})$ on $\mathcal{G}_{2}(G)$ is very often not transitive [Gel24, Remark 1.4]. Thus, there exists an epimorphism $F_{2}\to G$ and a generating pair for $G$ which cannot be lifted to $F_{2}$. However, the abelianisation map is trivial in this case. This demonstrates that the unique phenomenon exhibited in 1.3 for Lie groups, does not apply for general groups.
The Redundancy Rank.
Another natural invariant, which was studied before in many contexts, is the redundancy rank, which we denote by $\rho(G)$. It is the minimal $m\in\mathbb{N}\cup\left\{\infty\right\}$ such that, if $g_{1},\dots,g_{n}\in G$ generate $G$ and $n>m$, then one of the $g_{i}$’s is redundant (i.e., there is some $S\subsetneq\left\{g_{1},\dots,g_{n}\right\}$ that still generates $G$).
The redundancy rank has been extensively studied for finite groups, see [Whi00, LMS00, Luc13.1, Luc13.2, Har23] and [CC02, EG16, HHR07]. We are not aware of it ever having been given a name.
In Section 3, we show the following.
Theorem 1.9.
The redundancy rank of a connected abelian Lie group $A$ is equal to $\dim A+\dim A/T$, where $T$ is the maximal torus of $A$.
Using this and the results of [AN24], it is possible to reduce the computation of the redundancy rank of any connected Lie group to the computation of the redundancy rank of its maximal semisimple Lie subgroup. Computing the redundancy rank of these groups seems to be an intricate problem.
Question 1.10.
What is the redundancy rank of a connected simple Lie group?
Curiously, the redundancy rank of $\mathrm{SL}_{2}(\mathbb{R})$ and $\mathrm{SL}_{2}(\mathbb{C})$ is $\infty$! This was proved in [Min13] (although it is not explicitly stated there). Indeed, Minsky constructs for every $n\geq 2$, a generating $n$-tuple in $\mathrm{SL}_{2}(\mathbb{C})$ which is not redundant. In fact, he demonstrates a much stronger property, namely, that this $n$-tuple remains irredundant even after applying Nielsen transformations. We refer to [Gel24] for an excellent survey on actions of $\mathrm{Aut}(F_{n})$ on representation spaces, covering, among others, the result by Minsky.
It seems reasonable to conjecture that the redundancy rank of any non-compact semisimple Lie group is infinite. In a future work, we investigate the redundancy rank of connected compact simple Lie groups, as well as the Zariski redundancy rank of algebraic semisimple groups, which is analogously defined using the Zariski topology.
Paper Overview.
Section 2 is a preliminary section in which we introduce general tools and terminology useful for lifting generators. In Section 3 we restrict to the case $G$ is abelian. We show that every open epimorphism $f:G\to H$ is $n$-Gaschütz for every $n\geqslant\rho(G)$, and we show this bound is tight.
In Section 4 we consider semisimple Lie groups $G$. The main step there is to show the following.
Proposition 1.11.
Consider two connected simple Lie groups $G$ and $H$. Then any generating tuple $x_{1},...,x_{n}\in G$ can be paired with a (generating) tuple $y_{1},...,y_{n}\in H$ such that $(x_{1},y_{1}),...,(x_{n},y_{n})$ generate the product $G\times H$.
The key idea here is the use of suitable word maps to transfer elements of $G$ into a small ‘Zassenhaus neighbourhood’ of the identity. We believe the proof of this proposition may be useful for other generation problems in semisimple Lie groups.
In 5, we consider products of semisimple Lie groups with abelian Lie groups. In Section 6, we consider what we call Abels–Noskov groups, a very tame class of groups, the generating subsets of which were completely classified in [AN24] (see 6.3 and 6.4). Using their results, we can reduce the proof of 1.3 to Abels–Noskov groups in Section 7, thus concluding the proof. Finally, in Section 8 we discuss the Gaschütz rank of connected Lie groups.
Acknowledgement.
We extend our gratitude to Uri Bader, Tsachik Gelander, Alex Lubotzky and Gregory Soifer for inspiration and very helpful discussions. We thank Abels and Noskov for communications regarding this project as well as their closely related project [AN24] which we strongly rely on. Sadly, Noskov passed away before we finished the project.
Our research is partially supported by the ERC Advanced Grant 101052954 – SRS. The second author is supported by NSF postdoctoral fellowship DMS-2402368.
2 Preliminaries
In this section we establish our terminology and prove a few basic lemmas. Throughout this paper, we assume all homomorphisms are continuous; similarly, if $G$ is a topological group, we say that a subset $S\subseteq G$ generates $G$ if it topologically generates $G$: that is, if $\left\langle S\right\rangle$ (the group abstractly generated by $S$) is dense in $G$. Recall that $\mathrm{d}(G)$, the (topological) generating rank of $G$, is the minimal size of a generating set of $G$.
As we will see below, it is going to be useful for our purposes to consider non-Hausdorff groups. Accordingly, we do not assume topological groups are necessarily Hausdorff.
Notation 2.1.
If $f:G\to H$ is a map and $h_{1},\dots,h_{n}\in H$, we say that $g_{1},\dots,g_{n}\in G$ are lifts if $f(g_{i})=h_{i}$ for $i=1,\dots,n$.
Definition 2.2.
Let $G$ be a topological group, $f:G\to H$ an epimorphism. We say that $f$ is absolutely Gaschütz if, whenever $h_{1},\dots,h_{n}\in H$ generate $H$ (for arbitrary $n\in\mathbb{N}$), and $g_{1},\dots,g_{n}\in G$ are lifts, necessarily $g_{1},\dots,g_{n}$ generate $G$.
Definition 2.3.
Let $G$ be a topological group, $N\trianglelefteqslant G$. We say that $N$ is non-generating if, for every $H\leqslant G$ such that $HN$ is dense in $G$, necessarily $H$ is dense in $G$.
The following is immediate:
Lemma 2.4.
If $N$ is a non-generating normal subgroup of a topological group $G$, then $f:G\to G/N$ is absolutely Gaschütz.
In particular, if $N\trianglelefteqslant G$ is non-generating, then $\mathrm{d}(G)=\mathrm{d}(G/N)$. As for lifting generators, we have the following:
Lemma 2.5.
Let $G$ be a topological group, $f:G\to H$ an epimorphism. Let $N\trianglelefteqslant G$ be a non-generating normal subgroup, and consider the induced epimorphism $\tilde{f}:G/N\to H/f(N)$, which makes the following diagram commute:
[…]
Let $h_{1},\dots,h_{n}\in H$ be generators. Then they admit generating lifts to $G$ via $f$ if and only if $h_{1}f(N),\dots,h_{n}f(N)$ admit generating lifts to $G/N$ via $\tilde{f}$.
Observe that, even if $G$ and $H$ are Hausdorff, $f(N)$ might not be closed, so $H/f(N)$ might not be Hausdorff.
Proof.
Clearly, if $h_{1},\dots,h_{n}\in H$ admit generating lifts $g_{1},\dots,g_{n}\in G$, then $g_{1}N,\dots,g_{n}N$ are generating lifts of $h_{1}f(N),\dots,h_{n}f(N)$.
Now, in order to show the opposite direction, let $h_{1},\dots,h_{n}\in H$ be generators of $G$ such that $h_{1}f(N),\dots,h_{n}f(N)$ admit generating lifts $g_{1}N,\dots,g_{n}N\in G/N$. Since $G\to G/N$ is absolutely Gaschütz, we get that $g_{1},\dots,g_{n}$ generate $G$. It might seem that we are done, but not quite: $g_{1},\dots,g_{n}$ might not be lifts of $h_{1},\dots,h_{n}$. All we know is that $f(g_{i})f(N)=h_{i}f(N)$ for $i=1,\dots,n$. But this means that there are $\varepsilon_{1},\dots,\varepsilon_{n}\in N$ such that $f(g_{i}\varepsilon_{i})=h_{i}$ for $i=1,\dots,n$. Since $G\to G/N$ is absolutely Gaschütz, the elements $g_{1}\varepsilon_{1},\dots,g_{n}\varepsilon_{n}$ still generate $G$, so we are done. ∎
It will be useful to consider non-generating normal subgroups $N$ whose image in some quotients might not be closed; therefore, it is useful to consider quotients by non-closed normal subgroups as well, if we want to be able to quotient out any non-generating normal subgroup we like. Luckily, we don’t need to investigate quotients by any non-closed normal subgroup, such as $\mathbb{Q}$ in $\mathbb{R}$. Consider the diagram in the last lemma: while $\ker\tilde{f}$ might not be closed, it is still the image of $\ker f$ under the quotient $G\to G/N$, and $\ker f$ is closed. This gives us some information on $\ker f$, and hence $\ker\tilde{f}$. For instance, they are finitely generated:
Lemma 2.6.
A closed normal subgroup of a connected Lie group is (topologically) finitely generated.
Proof.
Let $N$ be a closed, normal subgroup of a connected Lie group $G$. Its connected component $N^{0}$ is a connected Lie group, hence finitely generated. Therefore, it is enough to show $N/N^{0}$ is finitely generated. Since $N^{0}$ is characteristic in $N$, it is normal in $G$, and $N/N^{0}$ is a normal subgroup of $G/N^{0}$. Thus, it is enough to show that discrete, normal subgroups of connected Lie group are finitely generated.
Let $\Gamma\trianglelefteqslant G$ be a discrete, normal subgroup of a connected Lie group $G$. Then $\Gamma$ is central in $G$. Consider the universal cover $\tilde{G}$ of $G$, and denote by $\pi:\tilde{G}\to G$ the covering map. Since $\pi$ is a local homeomorphism, we get that $\pi^{-1}(\Gamma)$ is discrete. It is also normal in $\tilde{G}$, hence central. Now, consider $\tilde{G}/\pi^{-1}(\Gamma)$; since $\tilde{G}$ is simply-connected, $\pi^{-1}(\Gamma)$ is the fundamental group of $\tilde{G}/\pi^{-1}(\Gamma)$. It is well-known that every connected Lie group is homotopy equivalent to a maximal compact subgroup, and that the fundamental group of a compact manifold is finitely generated. Since $\Gamma$ is a quotient of $\pi^{-1}(\Gamma)$, we get that $\Gamma$ is finitely generated. ∎
Observe that, if $\Delta\trianglelefteqslant G$ is a normal subgroup of a topological group $G$, then $g_{1},\dots,g_{n}\in G$ descend to generators of $G/\Delta$ if and only if they descend to generators of $G/\overline{\Delta}$, if and only if $\left\langle g_{1},\dots,g_{n}\right\rangle\Delta$ is dense.
It will sometimes be more convenient to focus on subgroups rather than maps. So, instead of considering an epimorphism $f:G\to H$ and looking for lifts of some generators $h_{1},\dots,h_{n}\in H$, we may begin with elements $g_{1},\dots,g_{n}\in G$ for which $\left\langle g_{1},\dots,g_{n}\right\rangle\ker f$ is dense in $G$ (i.e., elements descending to generators of $H$) and look for $k_{1},\dots,k_{n}$ such that $g_{1}k_{1},\dots,g_{n}k_{n}$ generate $G$.
3 Abelian Groups
If $A$ is a connected abelian Lie group, then its maximal compact subgroup $T$ is always isomorphic to a torus. Putting $m=\dim T$ and $n=\dim A-\dim T$, we always have $A\cong\mathbb{R}^{n}\times\mathbb{T}^{m}$.
Notation 3.1.
For the purposes of this section, the free rank of a connected abelian Lie group $A$ is $\dim A/T$, where $T$ is its maximal compact subgroup.
The purpose of this section is to show that, if $f:\mathbb{R}^{n}\times\mathbb{T}^{m}\to H$ is an open epimorphism and $h_{1},\dots,h_{k}\in H$ are generators of $H$ with $k\geqslant 2n+m$, there are lifts $g_{1},\dots,g_{k}\in\mathbb{R}^{n}\times\mathbb{T}^{m}$ that generate $\mathbb{R}^{n}\times\mathbb{T}^{m}$.
If $n=0$ and $H$ is Hausdorff, the requirement $k\geqslant 2n+m$ can be removed (e.g., by [CG18]); however, we will see that if either $n>0$ or $H$ is allowed not to be Hausdorff, there is an epimorphism $f:\mathbb{R}^{n}\times\mathbb{T}^{m}\to H$ and generators $h_{1},\dots,h_{2n+m-1}\in H$ of $H$ that do not admit generating lifts (see 3.10 and 3.12).
We begin with some preliminary lemmas.
Lemma 3.2.
A subset $S\subseteq\mathbb{R}^{n}\times\mathbb{T}^{m}$ generates $\mathbb{R}^{n}\times\mathbb{T}^{m}$ if and only if there is some $E\subseteq S$ projecting to a basis of $\mathbb{R}^{n}$ such that $S\backslash E$ projects to a generating set of $(\mathbb{R}^{n}\times\mathbb{T}^{m})/\langle E\rangle$. Moreover, $(\mathbb{R}^{n}\times\mathbb{T}^{m})/\langle E\rangle$ is isomorphic to $\mathbb{T}^{n+m}$.
Proof.
Denote by $p_{1}:\mathbb{R}^{n}\times\mathbb{T}^{m}\to\mathbb{R}^{n}$ the projection and set $G=\mathbb{R}^{n}\times\mathbb{T}^{m}$, $K=\mathbb{T}^{m}$. If $S\subseteq G$ generates $G$, then $p_{1}(S)$ generates $\mathbb{R}^{n}$, so clearly there is some $E\subseteq S$ such that $p_{1}(E)$ is a basis of $\mathbb{R}^{n}$, and obviously $S\backslash E$ projects to a generating set of $G/\langle E\rangle$ (since $S$ does, and $E$ projects to the trivial element there). Conversely, assume there is $E\subseteq S$ such that $p_{1}(E)$ is a basis of $\mathbb{R}^{n}$ and $S\backslash E$ projects to a generating set of $G/\langle E\rangle$. Denote $\tau:G\to G/\langle E\rangle$. Since $\tau(\left\langle S\right\rangle)$ is dense, we get that $\ker\tau\cdot\left\langle S\right\rangle$ is dense in $G$; but $\ker\tau\subseteq\left\langle S\right\rangle$, so we are done.
If $p_{1}(E)$ is a basis of $\mathbb{R}^{n}$, then, in particular, $p_{1}(\left\langle E\right\rangle)$ is closed; since $\ker p_{1}=K$ is compact, this means $\left\langle E\right\rangle$ is closed. Since it is countable and $G$ is locally compact, this means that it is discrete. Moreover, since $p_{1}(\left\langle E\right\rangle)$ is cocompact and $\ker p_{1}$ is compact, we get that $\left\langle E\right\rangle$ is cocompact; thus, $G/\left\langle E\right\rangle$ is a torus of dimension $n+m$. ∎
We will use the following standard lemma frequently.
Lemma 3.3.
Let $G$ be a connected abelian Lie group with maximal compact subgroup $K$, and let $\pi:G\to G/K$ be the projection. Let $g_{1},\dots,g_{n}\in G$. Then $g_{1},\dots,g_{n}$ generate a lattice in $G$ if and only if $\pi(g_{1}),\dots,\pi(g_{n})$ generate a lattice in $G/K$.
Proof.
If $\left\langle g_{1},\dots,g_{n}\right\rangle$ is a lattice then is it is particular closed, so $\left\langle g_{1},\dots,g_{n}\right\rangle K_{G}$ is closed, so $\pi(\left\langle g_{1},\dots,g_{n}\right\rangle)=\left\langle\pi(g_{1}),\dots,\pi(g_{n})\right\rangle$ is closed. Since it is countable and $G/K$ is locally compact, it is discrete. Since $\left\langle g_{1},\dots,g_{n}\right\rangle$ is cocompact, $\left\langle\pi(g_{1}),\dots,\pi(g_{n})\right\rangle$ is cocompact as well, hence a lattice.
Now, assume $\left\langle\pi(g_{1}),\dots,\pi(g_{n})\right\rangle$ is a lattice in $G/K$. Recall that $G/K$ is a vector group, and that this means $\pi(g_{1}),\dots,\pi(g_{n})$ are a basis. Denote $E=\{g_{1},\dots,g_{n}\}$; we claim $\left\langle E\right\rangle$ is discrete. Consider the homomorphism $\mathbb{Z}^{n}\longrightarrow\left\langle E\right\rangle$ sending the basis of $\mathbb{Z}^{n}$ to $E$, and the composition $\mathbb{Z}^{n}\longrightarrow\left\langle E\right\rangle\longrightarrow\pi(\left\langle E\right\rangle)$. It is onto, hence also injective (since $\mathbb{Z}^{n}$ is Hopfian); therefore, $\left\langle E\right\rangle$ does not intersect $K$. But if $\left\langle E\right\rangle$ is not discrete, then there is a sequence of nontrivial elements in $E$ approaching zero; since $\pi(\left\langle E\right\rangle)$ is discrete, almost all of them must be mapped to zero under $\pi$. This is a contradiction, since we just showed $\pi$ is injective on $E$. Since $K$ is compact and $\left\langle\pi(E)\right\rangle$ is cocompact, we get that $\left\langle E\right\rangle$ is discrete and cocompact, as needed. ∎
3.1 The Redundancy Rank
Definition 3.4.
A generating subset $S$ of a topological group $G$ is irredundant if, for every $s\in S$, the set $S\backslash\left\{s\right\}$ not longer generates $G$. The redundancy rank of a topological group $G$, denoted $\rho(G)$, is the maximal size of an irredundant finite generating subset. In other words, it is the minimal $m\in\mathbb{N}\cup\left\{\infty\right\}$ such that every generating subset $S$ with $\left|S\right|>n$ is redundant.
In this subsection we calculate the redundancy rank of connected abelian Lie groups.
Proposition 3.5.
The redundancy rank of $\mathbb{R}^{n}\times\mathbb{T}^{m}$ is $2n+m$ (for every $n,m\geqslant 0$).
Proof.
We need to prove two things: one, there is an irredundant generating set of $\mathbb{R}^{n}\times\mathbb{T}^{m}$ of size (at least) $2n+m$, and two, every finite generating set of $\mathbb{R}^{n}\times\mathbb{T}^{m}$ of size strictly greater than $2n+m$ has a redundant element.
First, by Kronecker’s theorem it’s easy to see that for any irrational $\alpha\in\mathbb{R}$ the sets $\{\alpha+\mathbb{Z}\}$ and $\{\alpha,1\}$ are irredundant generating sets of $\mathbb{T}$ and $\mathbb{R}$ respectively. Now, observe that if $S$ is an irredundant generating set of a group $G$ and $T$ is an irredundant generating set of a group $H$ then $S\times\{1_{H}\}\cup\{1_{G}\}\times T$ is an irredundant generating set of $G\times H$. Thus we can construct an irredundant generating set of $\mathbb{R}^{n}\times\mathbb{T}^{m}$ of size $2n+m$, as needed.
Now, let $S$ be a generating set of $\mathbb{R}^{n}\times\mathbb{T}^{m}$, $|S|>2n+m$. By Lemma 3.2, there is $E\subseteq S$ of size $n$ such that $S\backslash E$ projects to a generating set of $(\mathbb{R}^{n}\times\mathbb{T}^{m})/\langle E\rangle\cong\mathbb{T}^{n+m}$. So we just need to show that every generating set of a torus $G$ of size strictly greater than $\dim G$ has a redundant element. We prove this by induction on $\dim G$. Assume the claim is true for tori of dimension strictly smaller than $\dim G$, and let us prove it for $G$. Suppose $S\subseteq G$ generates $G$ and $|S|>\dim G$. Pick some $s_{0}\in S$, and let $n$ be the index of $\overline{\langle S\backslash\{s_{0}\}\rangle}^{0}$ inside $\overline{\langle S\backslash\{s_{0}\}\rangle}$. Then $T=\left\langle ns\right\rangle_{s\in S\backslash\{s_{0}\}}$ is of finite index in $\overline{\langle S\backslash\{s_{0}\}\rangle}$ and is contained in $\overline{\langle S\backslash\{s_{0}\}\rangle}^{0}$, hence it is equal to $\overline{\langle S\backslash\{s_{0}\}\rangle}^{0}$. If $T=G$, then $s_{0}$ is redundant and we are done; otherwise, $\dim T<\dim G$ and |$|\{ns\}_{s\in S\backslash\{s_{0}\}}|>\dim T,$ so, by the induction hypothesis, there is a redundant element. ∎
3.2 Lifting Generators
First, let us recall the following fact:
Lemma 3.6.
Let $G$ be a connected abelian Lie group, $A\leqslant G$ a closed connected subgroup. Then there is a closed connected subgroup $B$ such that the addition map $A\times B\to G$ is an isomorphism.
Lemma 3.7.
Let $G$ be an $n$-dimensional torus. Let $x_{1},\dots,x_{n}\in G$ be some elements, and let $\Delta\leqslant G$ be a (not-necessarily-closed) subgroup which is topologically finitely generated. Assume $\overline{\left\langle x_{i}\right\rangle}+\overline{\Delta}=G$. Then there are $\delta_{1},\dots,\delta_{n}\in\Delta$ such that $x_{1}+\delta_{1},\dots,x_{n}+\delta_{n}$ are topological generators of $G$.
Proof.
Set $A=\overline{\left\langle x_{i}\right\rangle_{i=1}^{n}}$, and set $k=[A:A^{0}]$. The group $\overline{\left\langle kx_{i}\right\rangle_{i=1}^{n}}$ is certainly contained in $A^{0}$, and is of finite index in it, and hence equals $A^{0}$. Set $m=\dim A$; the redundancy rank of $A\cong\mathbb{T}^{m}$ is $m$, so up to rearrangement we may assume $A^{0}=\overline{\left\langle kx_{i}\right\rangle_{i=1}^{m}}$.
Observe that $\overline{\Delta}^{0}\cong\mathbb{T}^{\dim\overline{\Delta}}$. Let $\delta_{1},\dots,\delta_{N}$ be topological generators of $\Delta$. Then they also topologically generate $\overline{\Delta}$. Therefore, setting $\ell=[\overline{\Delta}:\overline{\Delta}^{0}]$, we get that $\ell\delta_{1},\dots,\ell\delta_{N}$ topologically generate $\overline{\Delta}^{0}$. Thus, since the redundancy rank of $\overline{\Delta}^{0}$ is $\dim\overline{\Delta}$ and $\dim\overline{\Delta}\leqslant n$, there are some $n$ elements among $\ell\delta_{1},\dots,\ell\delta_{N}$ which already topologically generate $\overline{\Delta}^{0}$. Denote these elements (which are contained in $\Delta$) by $y_{1},\dots,y_{n}$, so that $\overline{\left\langle y_{i}\right\rangle_{i=1}^{n}}=\overline{\Delta}^{0}$.
Next, observe that, since $\overline{\Delta}^{0}$ is of finite index in $\overline{\Delta}$ and $A^{0}$ is of finite index in $A$, the group $A^{0}+\overline{\Delta}^{0}$ is of finite index in $A+\overline{\Delta}=G$. Since $G$ is connected, this means $G=A^{0}+\overline{\Delta}^{0}$, so $G$ is topologically generated by $\left\{kx_{i}\right\}_{i=1}^{n}\cup\left\{y_{j}\right\}_{j=1}^{n}$, and by the exact same argument one sees that $G$ is also topologically generated by $\left\{kx_{i}\right\}_{i=1}^{n}\cup\left\{ky_{j}\right\}_{j=1}^{n}$ (since once again the group topologically generated by this set is of finite index in $G$). Let $B$ be a closed connected subgroup of $G$ such that $G=A^{0}+B$ and $A^{0}\cap B=\left\{0\right\}$, and consider the well-defined projection $q:G\to B$. Since $\left\{kx_{i}\right\}_{i=1}^{n}\cup\left\{ky_{j}\right\}_{j=1}^{n}$ topologically generates $G$, its projection in $B$ topologically generates $B$; since $q(kx_{i})=0$ for each $i$, we see that $\left\{q(ky_{j})\right\}_{j=1}^{n}$ topologically generates $B$. Since the redundancy rank of $B\cong\mathbb{T}^{n-m}$ is $n-m$, we may assume (up to rearranging the $y_{j}$’s) that $B$ is already generated by $\left\{q(ky_{j})\right\}_{j=m+1}^{n}$. Since $\overline{\left\langle kx_{i}\right\rangle_{i=1}^{m}}=A^{0}$, it easily follows $\left\{kx_{i}\right\}_{i=1}^{m}\cup\left\{ky_{j}\right\}_{j=m+1}^{n}$ topologically generates $G$.
We are almost done. We now claim $\left\{x_{i}\right\}_{i=1}^{m}\cup\left\{x_{j}+y_{j}\right\}_{j=m+1}^{n}$ topologically generates $G$. This is immediate: the group topologically generated by this set obviously contains $kx_{i}$ for $i=1,\dots,m$, and hence contains $A^{0}$, and hence contains $kx_{j}$ for $j=m+1,\dots,n$ as well. Therefore, since it contains $x_{j}+y_{j}$ for $j=m+1,\dots,n$, it must also contain $ky_{j}$ for $j=m+1,\dots,n$. But we saw $\left\{kx_{i}\right\}_{i=1}^{m}\cup\left\{ky_{j}\right\}_{j=m+1}^{n}$ topologically generates $G$, so we are done. ∎
Lemma 3.8.
Let $G$ be a connected abelian Lie group, and let $\Delta$ be a (not-necessarily-closed) subgroup. Let $n$ and $\ell$ be the free ranks of $G$ and $G/\overline{\Delta}$ respectively. If $x_{1},\dots,x_{\ell}\in G$ are such that their projections in $G/\overline{\Delta}$ generate a lattice in $G/\overline{\Delta}$, and $x_{\ell+1},\dots,x_{n}\in G$ are any elements, then there are $\delta_{1},\dots,\delta_{n}\in\Delta$ such that $x_{1}+\delta_{1},\dots,x_{n}+\delta_{n}$ generate a lattice in $G$.
Proof.
Set $H=G/\overline{\Delta}$. Let $K_{G},K_{H}$ be the maximal tori of $G$ and $H$ respectively. Then $G/K_{G}\cong\mathbb{R}^{n}$ and $H/K_{H}\cong\mathbb{R}^{\ell}$. Since the image of $K_{G}$ in $H$ is contained in $K_{H}$, we have a well-defined epimorphism $\varphi:G/K_{G}\to H/K_{H}$, which is a linear map (as all continuous homomorphisms between vector groups over $\mathbb{R}$). Denoting by $\pi:G\to G/K_{G}$ and $p:H\to H/K_{H}$ the projections, we get the following commutative diagram:
${G\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${G/\Delta\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${G/K_{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${H/K_{H}}$
Denote by $\bar{x}_{1},\dots,\bar{x}_{n}$ the projections of $x_{1},\dots,x_{n}$ in $H$. Since $\bar{x}_{1},\dots,\bar{x}_{\ell}$ generate a lattice in $H$, their projections in $H/K_{H}$ generate a lattice in $H/K_{H}$. This means that $\left\{p(\bar{x}_{i})\right\}_{i=1}^{\ell}$ is a basis of $H/K_{H}$ when viewed as a vector space. Since $\varphi$ is linear, we obtain that $\pi(x_{1}),\dots,\pi(x_{\ell})\in G/K_{G}$ are linearly independent when we think of $G/K_{G}$ as a vector space (since they are lifts of $p(x_{1}),\dots,p(x_{\ell})$).
Next, we claim $\mathrm{Span}(\pi(\Delta))=\ker\varphi$. Clearly $\mathrm{Span}(\pi(\Delta))\subseteq\ker\varphi$. Now, suppose $\pi(g)\in\ker\varphi$; this means that the image of $g$ in $H$ is contained in $K_{H}$, which implies that for every identity neighbourhood $W$ of $H$ there is some nontrivial power of the image of $g$ which lies in it. Therefore, for every neighbourhood $W$ of $\overline{\Delta}$, there is some $n\geqslant 1$ such that $g^{n}\in W$. Clearly the same holds for neighbourhoods of $\Delta$. This means that for every neighbourhood $W$ of $\pi(\Delta)$ there is some $n\geqslant 1$ such that $n\pi(g)\in V$. Definitely the same holds for $\mathrm{Span}(\pi(\Delta))$, which is only bigger. But this obviously means that $ng\in\mathrm{Span}(\pi(\Delta))$ for some $n\geqslant 1$, so $\pi(g)\in\mathrm{Span}(\pi(\Delta))$, as needed.
Thus, we can extend the linearly independent set $\left\{\pi(x_{i})\right\}_{i=1}^{\ell}$ to a basis of $G/K_{G}$ using elements from $\pi(\Delta)$. In other words, there are $\varepsilon_{\ell+1},\dots,\varepsilon_{n}\in\Delta$ such that $\left\{\pi(x_{i})\right\}_{i=1}^{\ell}\cup\left\{\pi(\varepsilon_{j})\right\}_{j=\ell+1}^{n}$ is a vector space basis of $G/K_{G}$. Therefore, by the continuity of the determinant, there is some large $N\in\mathbb{N}$ such that $\left\{\pi(x_{i})\right\}_{i=1}^{\ell}\cup\left\{\pi(x_{j}+N\varepsilon_{j})\right\}_{j=\ell+1}^{n}$ is also a basis of $G/K_{G}$.
(To be more precise, consider the vector space decomposition $G/K_{G}=\mathrm{Sp}\left\{\pi(x_{i})\right\}_{i=1}^{\ell}\oplus\mathrm{Sp}\left\{\pi(\varepsilon_{j})\right\}_{j=\ell+1}^{n}$, which gives us a linear projection $G/K_{G}\to\mathrm{Sp}\left\{\pi(\varepsilon_{j})\right\}_{j=\ell+1}^{n}$. The projection of $\left\{\pi(\varepsilon_{j})\right\}_{j=\ell+1}^{n}$ in $\mathrm{Sp}\left\{\pi(\varepsilon_{j})\right\}_{j=\ell+1}^{n}$ is a basis, so, if $N\in\mathbb{N}$ is large enough, the projection of $\left\{\frac{1}{N}\pi(x_{j})+\pi(\varepsilon_{j})\right\}_{j=\ell+1}^{n}$ in $\mathrm{Sp}\left\{\pi(\varepsilon_{j})\right\}_{j=\ell+1}^{n}$ is still a basis of $\mathrm{Sp}\left\{\pi(\varepsilon_{j})\right\}_{j=\ell+1}^{n}$. Therefore $\left\{\pi(x_{i})\right\}_{i=1}^{\ell}\cup\left\{\frac{1}{N}\pi(x_{j})+\pi(\varepsilon_{j})\right\}_{j=\ell+1}^{n}$ is a basis of $G/K_{G}$, and hence $\left\{\pi(x_{i})\right\}_{i=1}^{\ell}\cup\left\{\pi(x_{j}+N\varepsilon_{j})\right\}_{j=\ell+1}^{n}$ is a basis as well.)
Denote $E=\{x_{1},\dots,x_{\ell},x_{\ell+1}+N\varepsilon_{\ell+1},\dots,x_{n}+N\varepsilon_{n}\}$, so that $\left\langle\pi(E)\right\rangle$ is a lattice of $G/K_{G}$, and hence $\left\langle E\right\rangle$ is a lattice of $G$ (by 3.3), and we are done (by setting $\delta_{i}=0$ for $i=1,\dots,\ell$ and $\delta_{i}=N\varepsilon_{i}$ for $i=\ell+1,\dots,n$). ∎
Proposition 3.9.
Let $G$ be a connected abelian Lie group with maximal torus $K$. If $\Delta$ is a (not-necessarily-closed) topologically-finitely-generated subgroup, $n\geqslant 2\dim G-\dim K$ and $g_{1},\dots,g_{n}\in G$ are such that $\left\langle g_{1},\dots,g_{n}\right\rangle\Delta$ is dense in $G$, then there are $\delta_{1},\dots,\delta_{n}\in\Delta$ such that $g_{1}+\delta_{1},\dots,g_{n}+\delta_{n}$ topologically generate $G$.
Proof.
Set $H=G/\overline{\Delta}$, and denote by $h_{1},\dots,h_{n}$ the images of $g_{1},\dots,g_{n}$, which are topological generators of $H$. By assumption, $n\geqslant\rho(G)$ (the redundancy rank of $G$).
Let $k$ and $\ell$ be the free ranks of $G$ and $H$ respectively, and let $K_{G},K_{H}$ be the tori of $G$ and $H$ respectively (so $G/K_{G}\cong\mathbb{R}^{k}$ and $H/K_{H}\cong\mathbb{R}^{\ell}$), and observe that $\ell\leqslant k$. Set $d=\dim G$, $m=\dim H$, so that $\rho(G)=d+k$ and $\rho(H)=m+\ell$.
Since $h_{1},\dots,h_{n}$ topologically generate $H$, their projections topologically generate $H/K_{H}$, and in particular span it as a vector space. Therefore, we may assume (up to rearrangement) that $h_{1},\dots,h_{\ell}$ project to a basis of $H/K_{H}$; this implies that their projections generate a lattice in $H/K_{H}$, and hence they generate a lattice in $H$. Now, since $h_{1},\dots,h_{n}$ topologically generate $H$ and $h_{1},\dots,h_{\ell}$ project to the trivial element in $H/\left\langle h_{i}\right\rangle_{i=1}^{\ell}$, we get that $h_{\ell+1},\dots,h_{d}$ project to topological generators of $H/\left\langle h_{i}\right\rangle_{i=1}^{\ell}\cong\mathbb{T}^{m}$. Since the redundancy rank of $H/\left\langle h_{i}\right\rangle_{i=1}^{\ell}$ is $m$, we can rearrange the $h_{i}$’s so that already $h_{n-m+1},\dots,h_{n}$ project to topological generators of $H/\left\langle h_{i}\right\rangle_{i=1}^{\ell}$. In particular, the $\ell+m$ elements $h_{1},\dots,h_{\ell},h_{n-m+1},\dots,h_{n}$ already topologically generate $H$. Observe that $n\geqslant d+k\geqslant m+\ell$, and hence there is no overlapping between the two sequences $1,\dots,\ell$ and $n-m+1,\dots,n$. In fact, since $n\geqslant d+k\geqslant m+k$, we even get that there is no overlapping between $1,\dots,k$ and $n-m+1,\dots,n$ (while $k\geqslant\ell$); we will use this below.
By 3.8, there are $\delta_{1},\dots,\delta_{k}\in\Delta$ such that $g_{1}+\delta_{1},\dots,g_{k}+\delta_{k}\in G$ generate a lattice in $G$. Now, clearly $g_{1},\dots,g_{\ell},g_{n-m+1},\dots,g_{n}$ together with $\Delta$ generate $G$ (since $h_{1},\dots,h_{\ell},h_{n-m+1},\dots,h_{n}$ generate $H$ and $\Delta$ is dense in $\ker(G\to H)$), so (since $g_{1},\dots,g_{k}$ project to nothing), $g_{n-m+1},\dots,g_{n}$ together with $\Delta$ project to generators of the torus $G/\left\langle g_{1},\dots,g_{k}\right\rangle$. Therefore 3.7 does the trick. ∎
3.3 Tightness
We now show that the bound given by 3.9 is tight in the case G is non-compact.
Proposition 3.10.
For every $n,m\geqslant 0$ there exists an open epimorphism $f:\mathbb{R}^{n+1}\times\mathbb{T}^{m}\to H$ onto a (Hausdorff) Lie group $H$ and a generating system $\underline{h}\in H^{2n+m+1}$ that doesn’t admit a generating lift.
We will need the following general version of Kronecker’s theorem:
Lemma 3.11 (Kronecker’s theorem).
For any index set $I$, the elements $(\alpha_{1}^{i}+\mathbb{Z})_{i\in I},\dots,(\alpha_{m}^{i}+\mathbb{Z})_{i\in I}$ in $\mathbb{T}^{I}$ generate $\mathbb{T}^{I}$ if and only if the following implication holds: if $\lambda_{1},\dots,\lambda_{n}\in\mathbb{Q}$ and $i_{1},\dots,i_{n}\in I$ are such that
$\lambda_{1}\alpha_{1}^{i_{1}}+\cdots+\lambda_{n}\alpha_{1}^{i_{n}}\in\mathbb{Q}$
$\vdots$
$\lambda_{1}\alpha_{m}^{i_{1}}+\cdots+\lambda_{n}\alpha_{m}^{i_{n}}\in\mathbb{Q}$
then $\lambda_{j}=0$ for every $j=1,\dots,n$.
Proof of 3.11.
This is immediate from Pontryagin duality. ∎
Proof of 3.10.
Let us write $G=(\mathbb{R}^{n}\times\mathbb{T}^{m})\times\mathbb{R}$ and set $H=(\mathbb{R}^{n}\times\mathbb{T}^{m})\times\mathbb{T}$, and consider the natural quotient $f:G\to H$. Set
[…]
$=\begin{pmatrix}\begin{matrix}1&&0\\ &\ddots\\ 0&&1\end{matrix}&\vline&\begin{matrix}\sqrt{2}&&0\\ &\ddots\\ 0&&\sqrt{2}\end{matrix}&\vline&0&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr 0&\vline&0&\vline&\begin{matrix}\sqrt{2}+\mathbb{Z}&&0\\ &\ddots\\ 0&&\sqrt{2}+\mathbb{Z}\end{matrix}&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr 0&\vline&0&\vline&0&\vline&\sqrt{2}+\mathbb{Z}\end{pmatrix}$
which is in $H^{2n+m+1}=H^{\rho(G)-1}$, and is a generating system of $H$, since $y_{1},\dots,y_{n+m},x$ project to generators of $H/\langle e_{i}\rangle_{i=1}^{n}$. Assume by contradiction there is a generating lift $\underline{g}=(\hat{e}_{1},\dots,\hat{e}_{n},\hat{y}_{1},\dots,\hat{y}_{n+m},\hat{x})\in G^{2n+m+1}$; it’s of the form
$\begin{pmatrix}\begin{matrix}1&&0\\ &\ddots\\ 0&&1\end{matrix}&\vline&\begin{matrix}\sqrt{2}&&0\\ &\ddots\\ 0&&\sqrt{2}\end{matrix}&\vline&0&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr 0&\vline&0&\vline&\begin{matrix}\sqrt{2}+\mathbb{Z}&&0\\ &\ddots\\ 0&&\sqrt{2}+\mathbb{Z}\end{matrix}&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr\begin{matrix}k_{1}&\cdots&k_{n}\end{matrix}&\vline&\begin{matrix}\ell_{1}&\cdots&\ell_{n}\end{matrix}&\vline&\begin{matrix}\ell_{n+1}&\cdots&\ell_{n+m}\end{matrix}&\vline&\sqrt{2}+k_{0}\end{pmatrix}$
for some $k_{0},\dots,k_{n},\ell_{1},\dots,\ell_{n+m}\in\mathbb{Z}$. Set $S=\{\hat{e}_{i}\}_{i=1}^{n}\cup\{\hat{x}\}$; since $\underline{g}$ generates $G$, we know in particular that $\{\hat{y}_{i}\}_{i=1}^{n+m}$ projects to a generating set of $G/\langle S\rangle$. We show this can’t be true.
Consider the universal covering $\tilde{G}=\mathbb{R}^{n+m+1}$, and lift $\underline{g}$ to $\underline{\tilde{g}}=(\tilde{e}_{1},\dots,\tilde{e}_{n},\tilde{y}_{1},\dots,\tilde{y}_{n+m},\tilde{x})\in\tilde{G}^{2n+m+1}$ thusly:
$\underline{\tilde{g}}=\begin{pmatrix}\begin{matrix}1&&0\\ &\ddots\\ 0&&1\end{matrix}&\vline&\begin{matrix}\sqrt{2}&&0\\ &\ddots\\ 0&&\sqrt{2}\end{matrix}&\vline&0&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr 0&\vline&0&\vline&\begin{matrix}\sqrt{2}&&0\\ &\ddots\\ 0&&\sqrt{2}\end{matrix}&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr\begin{matrix}k_{1}&\cdots&k_{n}\end{matrix}&\vline&\begin{matrix}\ell_{1}&\cdots&\ell_{n}\end{matrix}&\vline&\begin{matrix}\ell_{n+1}&\cdots&\ell_{n+m}\end{matrix}&\vline&\sqrt{2}+k_{0}\end{pmatrix}.$
Set $\varGamma=\langle\tilde{e}_{i}\rangle_{i=1}^{n}+\langle e_{n+i}\rangle_{j=1}^{m}+\langle\tilde{x}\rangle,$ so that $\tilde{G}/\varGamma=G/\langle S\rangle$ and $\{\tilde{y}_{i}\}_{i=1}^{n+m}$ projects to a generating set of $\tilde{G}/\varGamma$. From here we will arrive at a contradiction. Set
[…]
$=\begin{pmatrix}I_{n}&\vline&0&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr 0&\vline&I_{m}&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr\begin{matrix}k_{1}&\cdots&k_{n}\end{matrix}&\vline&\begin{matrix}0&\cdots&0\end{matrix}&\vline&\sqrt{2}+k_{0}\end{pmatrix}$
so that $\varPsi^{-1}$ sends $\varGamma$ to $\mathbb{Z}^{n+m+1}$, which means $\{\varPsi^{-1}\tilde{y}_{i}\}_{i=1}^{n+m}$ generates $\mathbb{R}^{n+m+1}/\mathbb{Z}^{n+m+1}$. One can see directly that
$\varPsi^{-1}=\begin{pmatrix}I_{n}&\vline&0&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr 0&\vline&I_{m}&\vline&\begin{matrix}0\\ \vdots\\ 0\end{matrix}\\ \hline\cr\begin{matrix}\frac{-k_{1}}{\sqrt{2}+k_{0}}&\cdots&\frac{-k_{n}}{\sqrt{2}+k_{0}}\end{matrix}&\vline&\begin{matrix}0&\cdots&0\end{matrix}&\vline&\frac{1}{\sqrt{2}+k_{0}}\end{pmatrix}$
so
$\varPsi^{-1}\tilde{y}_{i}=(0,\dots,\sqrt{2},\dots,0,\frac{\ell_{i}-k_{i}\sqrt{2}}{\sqrt{2}+k_{0}}),$
where $\sqrt{2}$ is in the $i$’th position and where we let $k_{i}\coloneqq 0$ for $i=n+1,\dots,m$. Therefore, by Kronecker’s theorem, there are no $\lambda_{1},\dots,\lambda_{n+m+1}\in\mathbb{Q}$ that aren’t all zero and $r_{1},\dots,r_{n+m}\in\mathbb{Q}$ such that, for all $i=1,\dots,n+m$,
$\lambda_{i}\sqrt{2}+\lambda_{n+m+1}\cdot\frac{\ell_{i}-k_{i}\sqrt{2}}{\sqrt{2}+k_{0}}+r_{i}=0.$ (1)
But by multiplying both sides by $\sqrt{2}+k_{0}$ we see that for each $i$ this equation is equivalent to
$\left(2\lambda_{i}+r_{i}k_{0}+\lambda_{n+m+1}\ell_{i}\right)+\left(\lambda_{i}k_{0}-\lambda_{n+m+1}k_{i}+r_{i}\right)\sqrt{2}=0,$
which is – since $\lambda_{i},r_{i},k_{i}$ are all rational – equivalent to the two equations
$\left(2\lambda_{i}+r_{i}k_{0}+\lambda_{n+m+1}\ell_{i}\right)$ $=0,$
$\left(\lambda_{i}k_{0}-\lambda_{n+m+1}k_{i}+r_{i}\right)$ $=0.$
These are $2n+2m$ homogeneous linear equations with $12n+2m+$ 1 variables ($\lambda_{1},\dots,\lambda_{n+m+1}$, $r_{1},\dots,r_{n+m}$), so there must be a nontrivial solution. But if one the $r_{i}$’s is nontrivial, one of the $\lambda_{i}$’s must be nontrivial as well (by Equation (1)). Contradiction! ∎
By [CG18], if $f:\mathbb{T}^{m}\to H$ is an epimorphism onto a Hausdorff group $H$, then any number of generators of $H$ can be lifted to $\mathbb{T}^{m}$. If we allow non-Hausdorff targets, the bound of 3.9 is tight.
Lemma 3.12.
For every $n\in\mathbb{N}$ there is an epimorphism $f:\mathbb{T}^{n}\to H$ and generators $h_{1},\dots,h_{n-1}\in H$ of $H$ that cannot be lifted to generators of $\mathbb{T}^{n}$.
Proof.
Let $x_{i}\in\mathbb{T}^{n}$ be the element whose $i^{\text{th}}$ coordinate is $\sqrt{2}$ and the rest of whose coordinates are zero. Set $\Delta=\left\langle x_{1},\dots,x_{n}\right\rangle$. Take $g_{i}=0$ for $i=1,\dots,n-1$. Then $\left\langle g_{1},\dots,g_{n-1}\right\rangle\Delta=\Delta$ is dense in $\mathbb{T}^{n}$, but there are no $\delta_{1},\dots,\delta_{n-1}\in\Delta$ such that $g_{1}+\delta_{1},\dots,g_{n-1}+\delta_{n-1}$ generate $\mathbb{T}^{n}$. ∎
4 Semisimple Groups
4.1 Preliminary Lemmas
Lemma 4.1.
Let $G$ be a topological group whose commutator is dense. If $F\leqslant G$ is a dense subgroup, then its commutator is also dense in $G$.
Proof.
Suppose $x\in G$ and $V\ni x$ is an open neighbourhood; then by assumption there are $x_{1},\dots,x_{2n}\in G$ such that $[x_{1},x_{2}]\cdots[x_{2n-1},x_{2n}]\in V$, which means there are open neighbourhoods $U_{i}$ of $x_{i}$ ($i=1,\dots,2n$) such that $[U_{1},U_{2}]\cdots[U_{2n-1},U_{2n}]\subseteq V$, and by assumption there are $f_{i}\in U_{i}\cap F$ ($i=1,\dots,2n$), so $[f_{1},f_{2}]\cdots[f_{2n-1},f_{2n}]$ belongs to $F^{\prime}$ and to $V$. ∎
Lemma 4.2.
Let $G$ be a topological group whose commutator is dense. Then every central subgroup of $G$ is non-generating; in particular, if $f:G\to H$ is an open homomorphism whose kernel is central, then $f$ is absolutely Gaschütz.
Proof.
If $Z\leqslant G$ is central and $F\leqslant G$ is a (not-necessarily-closed) subgroup such that $FZ$ is dense in $G$, then $[F,F]=[FZ,FZ]$ is dense in $[G,G]$, which is dense in $G$; in particular, $F\supseteq[F,F]$ is dense in $G$. In other words, every central subgroup is non-generating. ∎
We need a lemma from [GM17]. It is stated for groups, but the same proof works for Lie algebras as well.
Definition 4.3.
Let $\prod_{i=1}^{n}S_{i}$ be a product of Lie algebras, and denote by $p_{\ell}:\prod_{i=1}^{n}S_{i}\to S_{\ell}$ the projection. We say a subset $R\subseteq\prod_{i=1}^{n}S_{i}$ is diagonally embedded if there are two distinct $i,j\in\left\{1,\dots,n\right\}$ and an isomorphism $\varphi:S_{i}\to S_{j}$ such that $p_{j}(x)=\varphi(p_{i}(x))$ for all $x\in R$.
Lemma 4.4 ([GM17, Lemma 3.7]).
Let $S=\prod_{i=1}^{n}S_{i}$ be a product of simple Lie algebras and let $p_{i}:S\to S_{i}$ be the projection onto the $i^{\text{th}}$ coordinate. Let $R$ be a Lie subalgebra of $S$ that is not diagonally embedded and that satisfies $p_{i}(R)=S_{i}$ for every $i$. Then $R=S$.
We will use the following well known fact.
Fact ([Hum95, 1.5, 1.6]).
If $G$ is a connected simple Lie group and $g\in G$ is any element, then the closure of the orbit of $g$ under the group of (topological) automorphisms of $G$ is of measure zero.
Corollary 4.5.
Let $G=G_{1}\times\cdots\times G_{n}$ be a connected semisimple Lie group, $H$ a connected simple Lie group, $h\in H$ an element. Then
$\overline{\left\{(g_{i})_{i}\in G\middle|g_{i}=\varphi(h)\text{ for some $i$ and some isomorphism }\varphi:H\to G_{i}\right\}}$
is of measure zero.
4.2 Lifting Generators
In this section we show that connected semisimple Lie groups are Gaschütz.
For the sake of the proof, we call an identity neighbourhood $V$ in a Lie group $G$ a strongly Zassenhaus neighbourhood if $\log$ is a well defined diffeomorphism on it and if, whenever $(g_{1},g_{2})\in V\times V$ are such that $\log g_{1},\log g_{2}$ generate $\mathfrak{g}$, it follows $g_{1},g_{2}$ topologically generate $G$. Every connected semisimple Lie group admits a strongly Zassenhaus neighbourhood by [BG03, Theorem 2.1]. Moreover, by [BG03, Theorem 2.4] the subset of pairs $(g_{1},g_{2})\in V^{2}$ which do not topologically generate $G$ is contained in a closed null subset.
Lemma 4.6.
Let $G_{1},\dots,G_{m},H$ be connected centre-free (nontrivial) simple Lie groups, and let $h_{1},\dots,h_{n}\in H$ be topological generators of $H$ (in particular, $n\geq 2$). Then there are $g_{1},g_{2}\in G\coloneqq\prod_{i=1}^{m}G_{i}$ such that
$(g_{1},h_{1}),(g_{2},h_{2}),(1_{G},h_{3}),\dots,(1_{G},h_{n})\in G\times H$
topologically generate $G\times H$.
Proof.
Let $\Omega\subseteq H,W\subseteq G$ be strongly Zassenhaus neighbourhoods, and assume without loss of generality that $W\times\Omega$ is a strongly Zassenhaus neighbourhood in $G\times H$ as well.
Since the subset of pairs $(y_{1},y_{2})\in\Omega^{2}$ which do not generate $H$ is contained in a closed null subset, there are open subsets $V_{1},V_{2}\subseteq H$ such that every pair in $V_{1}\times V_{2}$ generates $H$. Since $h_{1},\dots,h_{n}$ generate $H$, there are words $w_{1},w_{2}$ in $n$ letters such that $w_{i}(\underline{h})\in V_{i}$ for $i=1,2$.
In fact, we can control the sum of powers of $h_{j}$ in $w_{i}(h_{1},\dots,h_{n})$ for every $j=1,\dots,n$ and $i=1,2$. To see this, think of words as elements in the free groups $F_{n}$ with basis $x_{1},\dots,x_{n}$. For $i=1,2$ and $j=1,\dots,n$, we can multiply $w_{i}$ from the right by $x_{j}\tilde{w}$, where $\tilde{w}\in[F_{n},F_{n}]$ is a word such that $\tilde{w}(\underline{h})\in x_{j}^{-1}w_{i}(\underline{h})^{-1}V_{i}$. This is possible since the commutator subgroup of $H$ is $H$, and therefore if $\varphi:F_{n}\to H$ is a homomorphism with a dense image, it follows $\varphi([F_{n},F_{n}])$ is dense as well (4.1). The new word (i.e. $w_{i}x_{j}\tilde{w}$) still satisfies the requirement above (i.e. $w_{i}x_{j}\tilde{w}(\underline{h})\in V_{i}$). By repeating this process, we can modify the sum of powers as we please. For reasons that will be clear momentarily, let us assume the following:
  1. 1.
    In $w_{1}$, the sum of powers of $x_{1}$ is $1$ and of $x_{2}$ is zero.
  2. 2.
    In $w_{2}$, the sum of powers of $x_{2}$ is $1$ and of $x_{1}$ is zero.
Now, consider the map
$\Psi:G\times G$ $\to G\times G$
$(g_{1},g_{2})$ $\mapsto(w_{1}(g_{1},g_{2},1_{G},\dots,1_{G}),w_{2}(g_{1},g_{2},1_{G},\dots,1_{G})).$
The differential of this map is
[…]
$(X_{1},X_{2})$ $\mapsto\left(n_{1}^{1}X_{1}+n_{2}^{1}X_{2},n_{1}^{2}X_{1}+n_{2}^{2}X_{2}\right),$
where $n_{j}^{i}$ is the sum of powers of $x_{j}$ in $w_{i}$. By our assumption, this implies that
$d\Psi(X_{1},X_{2})=(X_{1},X_{2}).$
In particular, the differential of $\Psi$ (at $(1_{G},1_{G})$) is of full rank, and hence the image of $\Psi$ contains an open neighbourhood of $(1_{G},1_{G})$. Let $U$ be the intersection of this identity neighbourhood with $W$, and let $U_{1},U_{2}\subseteq G$ be open subsets such that $U_{1}\times U_{2}\subseteq U$. As we remarked above, the subset $\mathcal{N}$ of $U_{1}\times U_{2}$ of pairs whose logarithms do not generate the Lie algebra of $G$ is contained a closed null subset.
Now, denote by $p_{\ell}:G\to G_{\ell}$ the projection onto the $\ell^{\text{th}}$ coordinate, for $\ell=1,\dots,m$. By 4.5, the closure of the subset
$\mathcal{B}\coloneqq\left\{(y_{1},y_{2})\in G^{2}|\begin{array}[]{l}\text{for some }\ell\text{ there is an isomorphism }\varphi:H\to G_{\ell}\\ \text{ such that }p_{\ell}(y_{i})=\varphi(w_{i}(\underline{h}))\text{ for }i=1,2\end{array}\right\}$
is null. Therefore we can find open subsets $\hat{U}_{i}\subseteq U_{i}$ ($i=1,2$) such that of
$\hat{U}_{1}\times\hat{U}_{2}\subseteq(U_{1}\times U_{2})\backslash\left(\mathcal{N}\cup\mathcal{B}\right).$
Let $g_{1},g_{2}\in G$ be such that $w_{i}(g_{1},g_{2},1_{G},\dots,1_{G})\in\hat{U}_{i}$ for $i=1,2$. Such an $n$-tuple actually exists since, by construction, the subset $\hat{U}_{1}\times\hat{U}_{2}$ lies in the image of $\Psi$. We claim $(g_{1},h_{1}),(g_{2},h_{2}),(1_{G},h_{3}),\dots,(1_{G},h_{n})\in G\times H$ generate $G\times H$. Set $\underline{g}=(g_{1},g_{2},1_{G},\dots,1_{G})$. Clearly, it is enough to show the pair
$\gamma_{1}\coloneqq(w_{1}(\underline{g}),w_{1}(\underline{h}))$ $=w_{1}((g_{1},h_{1}),(g_{2},h_{2}),(1_{G},h_{3}),\dots,(1_{G},h_{n})),$
$\gamma_{2}\coloneqq(w_{2}(\underline{g}),w_{2}(\underline{h}))$ $=w_{2}((g_{1},h_{1}),(g_{2},h_{2}),(1_{G},h_{3}),\dots,(1_{G},h_{n}))$
generates $G\times H$. The Lie algebra of $G\times H$ is $\mathfrak{g}\times\mathfrak{h}=\mathfrak{g}_{1}\times\cdots\times\mathfrak{g}_{n}\times\mathfrak{h}$; denote by $R$ the Lie subalgebra generated by $\log\gamma_{1},\log\gamma_{2}$. By our construction, $R$ projects onto $\mathfrak{h}$ and $\mathfrak{g}$, and therefore also onto each $\mathfrak{g}_{\ell}$ (for $\ell=1,\dots,m$). The subset $\left\{\log\gamma_{1},\log\gamma_{2}\right\}$ is not diagonally embedded: if it were, the isomorphism would be the differential of an isomorphism between the Lie groups (since the groups are connected, centre-free and semisimple), contrary to the construction of $\gamma_{1},\gamma_{2}$. Therefore, $R$ is not diagonally embedded, and hence $R=\mathfrak{g}\times\mathfrak{h}$ by 4.4. Since $\gamma_{1},\gamma_{2}\in W\times\Omega$ and $W\times\Omega$ is a strongly Zassenhaus neighbourhood, this means they generate $G\times H$. ∎
Corollary 4.7.
Connected semisimple Lie groups are Gaschütz. In other words, if $G$ is a connected semisimple Lie group, $f:G\to H$ is an open epimorphism and $h_{1},\dots,h_{n}\in H$ are generators with $n\geqslant\mathrm{d}(G)$, then there are generating lifts $g_{1},\dots,g_{n}\in G$.
Remark 4.8.
Recall that $\mathrm{d}(G)=2$ if $G$ is nontrivial, and that $H$ is necessarily a connected semisimple Lie group.
Proof.
For centre-free groups this follows immediately from the lemma, since connected centre-free semisimple Lie groups are direct products of connected centre-free simple Lie groups, and their only normal subgroups are subproducts, and hence every epimorphism from such a group is a composition of projections as in the lemma.
Now, let $G$ be a general semisimple Lie group, let $f:G\to H$ be an epimorphism and $\underline{h}\in H^{n}$ a generating system with $n\geqslant\mathrm{d}(G)$. Denote $K=\ker f$, $Z=Z(G)$, $\hat{G}=G/Z(G)$, $\hat{H}=H/f(Z)$. Observe $f(Z)\subseteq Z(H)$ is discrete and hence closed. We have a natural epimorphism $\hat{f}:\hat{G}\to\hat{H}$ and a commutative diagram
[…]
Clearly, $h_{1}f(Z),\dots,h_{n}f(Z)$ generate $\hat{H}$, so there are $g_{1}Z,\dots,g_{n}Z$ which generate $G/Z$ and such that $\hat{f}(g_{i}Z)=h_{i}f(Z)$. By 4.2, $g_{1},\dots,g_{n}$ generate $G$. It is still possible that they are not lifts of $h_{1},\dots,h_{n}$. However, the fact $\hat{f}(g_{i}Z)=h_{i}f(Z)$ means that $f(g_{i})^{-1}h_{i}\in f(Z)$ for $i=1,\dots,n$. Thus, there are $\varepsilon_{1},\dots,\varepsilon_{n}\in Z$ such that $f(\varepsilon_{i})=f(g_{i})^{-1}h_{i}$, so $f(g_{i}\varepsilon_{i})=h_{i}$ for $i=1,\dots,n$. Moreover, $g_{1}\varepsilon_{1},\dots,g_{n}\varepsilon_{n}$ generate $G$ (again by 4.2, because they are still lifts of $g_{1}Z,\dots,g_{n}Z$), so we are done. ∎
5 Reductive Groups
In this section we consider products of semisimple groups with abelian groups. Generation in such groups is simple:
Lemma 5.1 ([AN24, Lemma 6.6]).
Let $S$ be a connected semisimple Lie group and $A$ a connected abelian Lie group, and denote by $p_{1}:S\times A\to S$, $p_{2}:S\times A\to A$ the projections. Then $(s_{1},a_{1}),\dots,(s_{n},a_{n})\in S\times A$ generate $S\times A$ if and only if $s_{1},\dots,s_{n}\in S$ generate $S$ and $a_{1},\dots,a_{n}\in A$ generate $A$. In particular, $\mathrm{d}(S\times A)=\max\{\mathrm{d}(S),\mathrm{d}(A)\}$.
Proposition 5.2.
Let $S$ be a connected semisimple Lie group and $A$ a connected abelian Lie group. Let $f:S\times A\to H$ be an open epimorphism, and let $f^{\mathrm{ab}}:A\to H/f(S)$ be the induced epimorphism. If $h_{1},\dots,h_{n}\in H$ are generators with $n\geqslant\mathrm{d}(S\times A)$, then they admit generating lifts to $G$ via $f$ if and only if $h_{1}f(S),\dots,h_{n}f(S)$ admit generating lifts to $A$ via $f^{\mathrm{ab}}$.
If $S$ has finite centre and $H$ is Hausdorff, then $H/f(S)$ is Hausdorff as well.
Proof.
It is easy to see that, if $g_{1},\dots,g_{n}\in S\times A$ are generating lifts of $h_{1},\dots,h_{n}$, then their images in $A$ are generating lifts of $h_{1}f(S),\dots,h_{n}f(S)$. Thus, let $h_{1},\dots,h_{n}\in H$ be generators with $n\geqslant\mathrm{d}(S\times A)$ such that $h_{1}f(S),\dots,h_{n}f(S)$ admit generating lifts to $A$ via $f^{\mathrm{ab}}$. We will find generating lifts to $S\times A$.
Set $G=S\times A$ and let $\Delta$ be the kernel of $f$. Let $g_{1},\dots,g_{n}\in G$ be arbitrary lifts of $h_{1},\dots,h_{n}$. Since $h_{1},\dots,h_{n}$ generate $H$, we get that $\left\langle g_{1},\dots,g_{n}\right\rangle\Delta$ is dense in $G$; we need to prove there are $\delta_{1},\dots,\delta_{n}\in\Delta$ such that $g_{1}\delta_{1},\dots,g_{n}\delta_{n}$ generate $G$.
The case $\Delta\subseteq A$ or $\Delta\subseteq S$
If $\Delta\subseteq A$, then $(S\times A)/\Delta=S\times(A/\Delta)$. There are $s_{i}\in S,a_{i}\in A$ such that $h_{i}=(s_{i},a_{i}\Delta)$ for $i=1,\dots,n$. Since $a_{1}\Delta,\dots,a_{n}\Delta$ generate $A/\Delta$, there are $\varepsilon_{1},\dots,\varepsilon_{n}\in\Delta$ such that $a_{1}\varepsilon_{1},\dots,a_{n}\varepsilon_{n}$ generate $A$ (since the map $A\longrightarrow A/\Delta$ is exactly $f^{\mathrm{ab}}$), so $(s_{1},a_{1}\varepsilon_{1}),\dots,(s_{n},a_{n}\varepsilon_{n})$ are lifts of $(s_{1},a_{1}\Delta),\dots,(s_{n},a_{n}\Delta)$ that generate $S\times A$ (by 5.1), as needed.
The case $\Delta\subseteq S$ is exactly the same, only naturally one needs to use the fact $n\geqslant\mathrm{d}(S)$ and 4.7.
The case $\Delta$ is discrete
The centre of $S$ is non-generating as a subgroup of $S$ (4.2), and hence as a subgroup of $S\times A$ (by 5.1). Denote it by $Z$; we have the following commutative diagram:
[…]
where $\tilde{f}:(S/Z\times A)\to H/f(Z)$ is the epimorphism induced by $f$. By 2.5, the elements $h_{1},\dots,h_{n}\in H$ admit generating lifts to $S\times A$ if and only if the images of $h_{1},\dots,h_{n}$ in $H/f(Z)$ admit generating lifts to $(S/Z\times A)$ via $\tilde{f}$, so we may have assumed $S$ is centre-free. If $S$ is centre-free and $\Delta$ is discrete, $\Delta$ is contained in $A$ (being discrete and normal, hence central), so we may proceed as above.
The general case
Set $N=\overline{\Delta}$ and consider $N^{0}$. Since $\left\langle g_{1},\dots,g_{n}\right\rangle\Delta$ is dense in $G$, we get that $g_{1}N,\dots,g_{n}N$ generate $G/N$, which (by the isomorphism $G/N\cong(G/N^{0})/(N/N^{0})$) is the same thing as saying $(g_{1}N^{0})N/N^{0},\dots,(g_{n}N^{0})N/N^{0}$ generate $(G/N^{0})/(N/N^{0})$. Since $G/N^{0}$ is still a direct product of a semisimple Lie group and an abelian Lie group, and $N/N^{0}$ is a closed normal subgroup, we may proceed as in the case $\Delta$ is discrete. This means that there are $\varepsilon_{1}N^{0},\dots,\varepsilon_{n}N^{0}\in N/N^{0}$ such that $g_{1}\varepsilon_{1}N^{0},\dots,g_{n}\varepsilon_{n}N^{0}$ generate $G/N^{0}$. Now, $\Delta$ is dense in $N$, hence its image in $N/N^{0}$ is dense in $N/N^{0}$; but $N/N^{0}$ is discrete, so it means $\Delta$ surjects onto it, so we may assume $\varepsilon_{1},\dots,\varepsilon_{n}\in\Delta$. Thus, $g_{1}\varepsilon_{1}N^{0},\dots,g_{n}\varepsilon_{n}N^{0}\in G/N^{0}$ generate $G/N^{0}$, and we need to prove there are $\delta_{1},\dots,\delta_{n}\in\Delta$ such that $g_{1}\varepsilon_{1}\delta_{1},\dots,g_{n}\varepsilon_{n}\delta_{n}$ generate $G$.
In order to simplify notations, let us just assume $N$ is connected. So we are now in the following scenario: $G=S\times A$ where $S$ is a connected semisimple Lie group and $A$ is a connected abelian Lie group, $\Delta\trianglelefteqslant G$ is a not-necessarily-closed normal subgroup such that $N=\overline{\Delta}$ is connected, and $g_{1},\dots,g_{n}\in G$ are elements such that $g_{1}N,\dots,g_{n}N$ generate $G/N$. We need to prove there are $\delta_{1},\dots,\delta_{n}\in\Delta$ such that $g_{1}\delta_{1},\dots,g_{n}\delta_{n}$ generate $G$.
Denote by $p_{S}:S\times A\to S$ the projection, and consider $p_{S}(\Delta)$. This is a normal subgroup of $S$, hence a closed subgroup (see for example [Rag72]). We have $p_{S}(N)=p_{S}(\overline{\Delta})\subseteq\overline{p_{S}(\Delta)}=p_{S}(\Delta)$, and clearly $p_{S}(\Delta)\subseteq p_{S}(N)$, so $p_{S}(N)=p_{S}(\Delta)$. Denote $N_{S}=p_{S}(N)=p_{S}(\Delta)$. This is a connected normal subgroup of a semisimple Lie group, hence also a connected semisimple Lie group, which means in particular that $[N_{S},N_{S}]=N_{S}$. Now, consider $[\Delta,\Delta]$; on the one hand,
$p_{S}([\Delta,\Delta])=[p_{S}(\Delta),p_{S}(\Delta)]=[N_{S},N_{S}]=N_{S}.$
On the other hand, if $p_{A}:S\times A\to A$ is the projection onto $A$, we clearly have that $p_{A}([\Delta,\Delta])\leqslant A$ is trivial. In other words, $[\Delta,\Delta]$ is contained in $S$, which means that $[\Delta,\Delta]=N_{S}$ (so, in particular, it is closed). Clearly $[\Delta,\Delta]\subseteq\Delta$, so $N_{S}=p_{S}(\Delta)\subseteq\Delta$. That is, $\Delta\trianglelefteqslant S\times A$ contains its own projection into $S$, so $\Delta=p_{S}(\Delta)\times p_{A}(\Delta)=N_{S}\times p_{A}(\Delta)$. In particular, this means that $N=N_{S}\times\overline{p_{A}(\Delta)}$.
The rest follows from the case $\Delta$ is contained in either $A$ or $S$, since dividing by $N_{S}\times\rho_{A}(\Delta)$ is the same thing as first dividing by $N_{S}$ and then by (the image of) $\rho_{A}(\Delta)$.
Observe that, if $\Delta=N_{S}\times p_{A}(\Delta)$ is closed, $p_{A}(\Delta)$ must be closed as well, so $\ker f^{\mathrm{ab}}=p_{A}(\Delta)$ is closed. For this part, we need to assume the centre of $S$ is finite, because otherwise we might have changed the closedness of $\Delta\cdot S$ when dividing by it above. ∎
Remark 5.3.
If $S$ has infinite centre, it is possible for $H/f(S)$ not to be Hausdorff even though $H$ is; see 8.2.
6 Abels–Noskov Groups
At this point, we have handled the cases where $G$ is semisimple, abelian, or a product of such groups. Extending this to the general case involves using the structure theory of Lie groups. Fortunately, we can rely on the work of [AN24] where the Frattini subgroup of a connected Lie group is characterised. Elements of this subgroup are irrelevant to generation problems, so by quotienting out this subgroup, we obtain a group with a highly structured form, which we refer to as Abels–Noskov groups.
This section is dedicated to establishing Theorem 1.3 for this specific class. In the following section 7, we will show how the proof of the general case (Theorem 1.3) can be reduced to this setting.
6.1 Generating Sets of Modules
We start by collecting facts about finite dimensional representations of abstract groups, and their generating sets.
Let $L$ be a group, $V$ a finite dimensional vector space over $\mathbb{R}$ and $\rho:L\to\mathrm{GL}(V)$ a representation. We denote by $\mathbb{R}[\rho]$ the subalgebra of $\mathrm{End}_{\mathbb{R}}(V)$ generated by $\rho(L)$. We call $V$ an $L$-module, and we call a subspace $U$ of $V$ a submodule if it is $L$-invariant (equivalently, $\mathbb{R}[\rho]$-invariant). In this case, we denote by $\rho|_{U}:L\to\mathrm{GL}(U)$ the corresponding representation. We say $V$ (or $\rho$) is irreducible if it admits no nontrivial submodules. It is basic fact in representation theory that $V$ is a direct sum of irreducible submodules if, and only if, every submodule admits a complement (i.e., every submodule $U\subseteq V$ admits a submodule $W\subseteq V$ such that $V=U\oplus W$ as vector spaces, and hence as $L$-modules). In this case we say that $V$ is completely reducible. For a subset $S\subseteq V$, the submodule generated by $S$ is the smallest submodule of $V$ containing $S$.
We say that $V$ is isotypic if it is a direct sum of irreducible representations which are all isomorphic to one another. The decomposition is not unique, but obviously the number of irreducible submodules in the sum is; it is called the multiplicity of the isotypic module $V$. We denote it by $m(V)$.
Every completely reducible module is a direct sum of isotypic submodules, and this decomposition is unique. Given an irreducible representation $\sigma$ of $L$, we denote by $V_{\sigma}$ the isotypic component of $V$ corresponding to $\sigma$ (which might be zero). The $\sigma$-multiplicity of $V$ is the multiplicity of $V_{\sigma}$. We denote it by $m_{\sigma}(V)$. We denote $m(V)=\max_{\sigma}\left\{m_{\sigma}(V)\right\}$.
If $\sigma:L\to\mathrm{GL}(V)$ is an irreducible representation, then the centraliser of $\mathbb{R}[\sigma]$ inside $\mathrm{End}_{\mathbb{R}}(V)$ is a division algebra $k$ over $\mathbb{R}$, and as such $k$ is isomorphic to $\mathbb{R}$, $\mathbb{C}$ or $\mathbb{H}$ (the quaternions). In this case we may regard $V$ as a vector space over $k$, and it holds that $\mathbb{R}[\sigma]\cong\mathrm{End}_{k}(V)$. If $\rho$ is isotypic of type $\sigma$, then the diagonal embedding gives us an isomorphism $\mathbb{R}[\sigma]\to\mathbb{R}[\rho]$.
Given a representation $\rho:L\to\mathrm{GL}(V)$, we get another module, namely $\mathbb{R}[\rho]$ itself, where the action of $L$ on $\mathbb{R}[\rho]$ is given by .$\ell.A=\rho(\ell)\circ A\in\mathbb{R}[\rho]$. If $\rho$ is irreducible, then $\mathbb{R}[\rho]$ is isotypic of type $\rho$, and its multiplicity is $\dim_{k}V$.
Lemma 6.1 ([AN24, Corollary C.2]).
Let $\sigma$ be an irreducible representation of a group $L$, and let $\rho:L\to\mathrm{GL}(V)$ be an isotypic representation of $L$ of type $\sigma$ on a finite dimensional real vector space $V$, which is isomorphic to the sum of $\ell$ copies of $\sigma$. Denote by $\mathrm{d}_{\mathbb{R}[\rho]}(V)$ the minimal number of elements needed to generate $V$ as an $\mathbb{R}[\rho]$-module. Using the notations above, we have
$\mathrm{d}_{\mathbb{R}[\rho]}(V)=\left\lceil\frac{\ell}{\dim_{k}V}\right\rceil=\left\lceil\frac{\dim_{\mathbb{R}}V}{\dim_{\mathbb{R}}\mathbb{R}[\rho]}\right\rceil.$
Corollary 6.2 ([AN24, Corollary C.3]).
Let $\rho:L\to\mathrm{GL}(V)$ be a completely reducible finite dimensional representation of a group $L$ on real vector space $\mathbb{R}$ with isotypic components $\left\{V_{\sigma}\right\}_{\sigma}$. Let $k_{\sigma}$ be the centraliser of $\mathbb{R}[\rho|_{V_{\sigma}}]$, so that $V_{\sigma}$ is a vector space over the division algebra $k_{\sigma}$. Denote by $\mathrm{d}_{\mathbb{R}[\rho]}(V)$ the minimal number of elements needed to generate $V$ as an $\mathbb{R}[\rho]$-module. Using the notations above, we have
$\mathrm{d}_{\mathbb{R}[\rho]}(V)=\max_{\sigma}\left\{\left\lceil\frac{m_{\sigma}(V)}{\dim_{k_{\sigma}}V_{\sigma}}\right\rceil\right\}=\max_{\sigma}\mathrm{d}_{\mathbb{R}[\rho|_{V_{\sigma}}]}(V_{\sigma}).$
6.2 Generating Sets of Abels–Noskov Groups
Definition 6.3.
An Abels–Noskov group is a connected Lie group of the form $(S\times A)\ltimes_{\rho}V$, where
  1. 1.
    $S$ is a connected semisimple Lie group with finite centre;
  2. 2.
    $A$ is a connected abelian Lie group;
  3. 3.
    $V$ is a finite dimensional real vector space;
  4. 4.
    The semi-direct product is defined via a representation $\rho:(S\times A)\to\mathrm{GL}(V)$ that is completely reducible and admits no non-trivial fixed vectors.
We may now quote the beautiful result of Abels and Noskov which characterizes generating sets.
Theorem 6.4 ([AN24, Lemmas 6.4 and 6.5]).
Let $L\ltimes_{\rho}V$ be an Abels–Noskov group and let $(\ell_{1},v_{1}),\dots,(\ell_{n},v_{n})\in L\ltimes_{\rho}V$. Let $\varepsilon_{\boldsymbol{\ell}}:\mathbb{R}[\rho]^{n}\to\mathbb{R}[\rho]$ be the $\mathbb{R}[\rho]$-module homomorphism
$\varepsilon_{\boldsymbol{\ell}}(a_{1},\dots,a_{n})=\sum_{i=1}^{n}a_{i}(\boldsymbol{1}-\rho(\ell_{i}))$
and let $\alpha_{\boldsymbol{v}}:\mathbb{R}[\rho]^{n}\to V$ be the $\mathbb{R}[\rho]$-module homomorphism
$\alpha_{\boldsymbol{v}}(a_{1},\dots,a_{n})=\sum_{i=1}^{n}a_{i}v_{i}.$
Then $(\ell_{1},v_{1}),\dots,(\ell_{n},v_{n})$ generate $L\ltimes_{\rho}V$ if and only if the following two conditions hold: $\ell_{1},\dots,\ell_{n}$ generate $L$, and
$\alpha_{\boldsymbol{v}}(\ker\varepsilon_{\boldsymbol{\ell}})=V.$
Observe that every $\mathbb{R}[\rho]$-module homomorphism $\mathbb{R}[\rho]^{n}\to\mathbb{R}[\rho]$ is of the form $\varepsilon_{\boldsymbol{\ell}}$ for $\boldsymbol{\ell}=(\varepsilon(e_{1}),\dots,\varepsilon(e_{n}))$ and every $\mathbb{R}[\rho]$-module homomorphism $\mathbb{R}[\rho]^{n}\to V$ is of the form $\alpha_{\boldsymbol{v}}$ for $\boldsymbol{v}=(\alpha(e_{1}),\dots,\alpha(e_{n}))$, where $e_{i}\in\mathbb{R}[\rho]^{n}$ is $\boldsymbol{1}$ in the $i^{\text{th}}$ coordinate and 00 in the others. If $\ell_{1},\dots,\ell_{n}$ generate $L$, then $\varepsilon_{\boldsymbol{\ell}}$ is always surjective, by [AN24, Lemma C.4].
6.3 The Case $\ker f\subseteq V$
Lemma 6.5.
Let $G=L\ltimes_{\rho}V$ be an Abels–Noskov group, let $\left\{V_{\sigma}\right\}_{\sigma}$ be the isotypic components of $V_{\sigma}$, and let $\pi_{\sigma}:L\ltimes_{\rho}V\to L\ltimes_{\rho|_{V_{\sigma}}}V_{\sigma}$ be the projections. Let $X\subseteq G$. Then $X$ generates $G$ if and only if $\pi_{\sigma}(X)$ generates $L\ltimes_{\rho|_{V_{\sigma}}}V_{\sigma}$ for every $\sigma$.
Proof.
First, recall that if $U\subseteq V$ is a submodule, then the isotypic components of $U$ are $U\cap V_{\sigma}$, and that $U\cap V_{\sigma}=p_{\sigma}(U)$ (where $p_{\sigma}:V\to V_{\sigma}$ is the projection). In particular, if $U$ projects onto every isotypic component $V_{\sigma}$, then $U=V$. With this in mind, the lemma follows immediately from 6.4. ∎
Lemma 6.6.
Let $G=L\ltimes_{\rho}V$ be an Abels–Noskov group, and let $f:G\to H$ be an open epimorphism such that $\ker f$ is contained in $V$. If $h_{1},\dots,h_{n}\in H$ generate $H$ and $n\geqslant\mathrm{d}(G)$ then there are lifts $g_{1},\dots,g_{n}\in G$ that generate $G$.
Proof.
Write $U=\ker f$; since it is a normal subgroup of $G$ contained in $V$, it is an $L$-submodule of $V$ (see [AN24, Lemma 6.4]). [^5] In particular, even if we don’t assume it to be closed, it actually must be, since vector subspaces are always closed (in finite dimension). Therefore, there is an $L$-submodule $W\subseteq V$ such that $V=W\oplus U$, and $H$ is naturally isomorphic to $L\ltimes_{\rho|_{W}}W$.
In light of 6.5, we may assume $\rho$ is an isotypic representation. As usual, we denote by $\mathbb{R}[\rho]$ the subalgebra of $\mathrm{End}(V)$ generated by $\rho(L)$. It is isomorphic to $\mathbb{R}[\rho|_{W}]$, the subalgebra of $\mathrm{End}(W)$ generated by $\rho|_{W}(L)$ (except in the case $U=V$, in which case everything below works with minor changes). Writing $h_{i}=(\ell_{i},w_{i})$ for $i=1,\dots,n$, we denote by $\varepsilon_{\boldsymbol{\ell}}:\mathbb{R}[\rho]^{n}\to\mathbb{R}[\rho]$ the map
$\varepsilon_{\boldsymbol{\ell}}\left(a_{1},\dots,a_{n}\right)=\sum_{i=1}^{n}a_{i}(\boldsymbol{1}-\rho(\ell_{i})),$
as above. Abusing notations somewhat, we identify $\mathbb{R}[\rho]$ with $\mathbb{R}[\rho|_{W}]$ and denote by $\varepsilon_{\boldsymbol{\ell}}$ also the map $\mathbb{R}[\rho|_{W}]^{n}\to\mathbb{R}[\rho|_{W}]$ defined by
$\varepsilon_{\boldsymbol{\ell}}\left(a_{1},\dots,a_{n}\right)=\sum_{i=1}^{n}a_{i}(\boldsymbol{1}-\rho|_{W}(\ell_{i})).$
We write $P=\ker\varepsilon_{\boldsymbol{\ell}}$. We denote by $\alpha_{\boldsymbol{w}}:\mathbb{R}[\rho]^{n}\to W$ the map
$\alpha_{\boldsymbol{w}}\left(a_{1},\dots,a_{n}\right)=\sum_{i=1}^{n}a_{i}w_{i}.$
The fact $h_{1},\dots,h_{n}$ generate $H$ means that $\alpha_{\boldsymbol{w}}(P)=W$.
We need to find lifts $g_{1},\dots,g_{n}\in G$ that generate $G$. This means finding $u_{1},\dots,u_{n}\in U$ such that $(\ell_{1},w_{1}\oplus u_{1}),\dots,(\ell_{n},w_{n}\oplus u_{n})$ generate $G$. For $\boldsymbol{u}=(u_{1},\dots,u_{n})\in U^{n}$, we denote by $\alpha_{\boldsymbol{u}}:\mathbb{R}[\rho]^{n}\to U$ the map
$\alpha_{\boldsymbol{u}}\left(a_{1},\dots,a_{n}\right)=\sum_{i=1}^{n}a_{i}u_{i}.$
We denote by $\alpha_{\boldsymbol{w}\oplus\boldsymbol{u}}:\mathbb{R}[\rho]^{n}\to V$ the map
$\alpha_{\boldsymbol{w}\oplus\boldsymbol{u}}\left(a_{1},\dots,a_{n}\right)=\sum_{i=1}^{n}a_{i}(w_{i}\oplus u_{i}),$
so that $\alpha_{\boldsymbol{w}\oplus\boldsymbol{u}}=\alpha_{\boldsymbol{w}}\oplus\alpha_{\boldsymbol{u}}$. Therefore, we need to find $\boldsymbol{u}=(u_{1},\dots,u_{n})\in U^{n}$ such that $\left(\alpha_{\boldsymbol{w}}\oplus\alpha_{\boldsymbol{u}}\right)(P)=V$. As noted above, every map $\alpha:\mathbb{R}[\rho]^{n}\to U$ is of the form $\alpha_{\boldsymbol{u}}$ for some $\boldsymbol{u}=(u_{1},\dots,u_{n})\in U^{n}$ (namely, $\boldsymbol{u}=(\alpha(e_{1}),\dots,\alpha(e_{n}))$), so we actually just need to find a map $\alpha:\mathbb{R}[\rho]^{n}\to U$ such that $\left(\alpha_{\boldsymbol{w}}\oplus\alpha\right)(P)=V$. Since $\mathbb{R}[\rho]^{n}\cong P\oplus\mathbb{R}[\rho]$, this is the same thing as finding a map $\alpha:P\to U$ such that $\left(\bar{\alpha}_{\boldsymbol{w}}\oplus\alpha\right)(P)=V$, where $\bar{\alpha}_{\boldsymbol{w}}$ is the restriction of $\alpha_{\boldsymbol{w}}$ to $P$.
Write $Q=\ker\bar{\alpha}_{\boldsymbol{w}}$. Since $P$ is completely reducible, we get that $P\cong Q\oplus W$. Observe that there is a surjective $\mathbb{R}[\rho]$-map $\beta:Q\to U$ if (and only if) $m(Q)\geqslant m(U)$. Recall that
$m(\mathbb{R}[\rho])=\dim_{k}V,$
where $k$ is the corresponding Schur field (i.e., the centraliser of $\mathbb{R}[\rho]$). Thus,
$m(P)=(n-1)\cdot m(\mathbb{R}[\rho])=(n-1)\dim_{k}V.$
So we know that
$m(Q)=m(P)-m(W)=(n-1)\dim_{k}V-m(W).$
So what we need is that
$(n-1)\dim_{k}V\geqslant m(W)+m(U)=m(V).$
That is, we need
$n\geqslant\frac{m(V)}{\dim_{k}V}+1.$
As luck would have it, our assumption on $n$ is that
$n\geqslant\mathrm{d}(G)\geqslant d_{\mathbb{R}[\rho]}(V)+1=\left\lceil\frac{m(V)}{\dim_{k}V}\right\rceil+1,$
so we’re in the clear. So there is such a $\beta$, and we can define $\alpha:Q\oplus W\to U$ by $\alpha(q,w)=\beta(q)$. Observe that, with the identification $P\cong Q\oplus W$, the map $\bar{\alpha}_{\boldsymbol{w}}:P\to W$ is the projection onto the second coordinate (it was, after all, exactly the map $\bar{\alpha}_{\boldsymbol{w}}$ that gave us this isomorphism). Therefore, $\left(\bar{\alpha}_{\boldsymbol{w}}\oplus\alpha\right)(W)=W$ and $\left(\bar{\alpha}_{\boldsymbol{w}}\oplus\alpha\right)(Q)=U$, so $\left(\bar{\alpha}_{\boldsymbol{w}}\oplus\alpha\right)(P)=V$, as needed. ∎
6.4 The Case $\ker f\cap V$ is Trivial
Lemma 6.7.
Let $G=L\ltimes_{\rho}V$ be an Abels–Noskov group, and let $K\trianglelefteqslant G$ be a not-necessarily-closed normal subgroup of $G$ such that $K\cap V$ is trivial. Then $K\leqslant\ker\rho\leqslant L$
Proof.
Let $\pi:G\to L$ denote the quotient map. Since $V$ is abelian, we have, for any $k\in K$ and $v\in V$,
$[k,v]=[\pi(k),v]=\rho(\pi(k))v-v.$
Moreover, $[k,v]\in K\cap V$, since both $K$ and $V$ are normal, and $K\cap V$ is trivial. Therefore, $K$ and $\pi(K)$ act trivially on $V$ by conjugation; since $\pi(K)$ is contained in $L$ (and the action by conjugation is given by $\rho$), we have that $\pi(K)\leqslant\ker\rho$.
As $K$ is normal in $G$, $\overline{\pi(K)}^{0}$ is normal in $L$. Any connected simple normal subgroup of $S$ is either contained in $\overline{\pi(K)}^{0}$ or else its intersection with it is trivial. Let $H_{0}$ be the product all connected simple normal subgroups of $S$ which are not contained in $\overline{\pi(K)}^{0}$, and set $H=H_{0}A$. Then $\overline{\pi(K)}^{0}H=L$ and $[\overline{\pi(K)}^{0},H]=\{1\}$. Now, let $k\in K$. There is some $v_{k}\in V$ such that $k=\pi(k)v_{k}$. We have seen that $\pi(K)$ commutes with $H$ as well as with $V$. Hence for any $h\in H$ we have that
$[k,h]=[\pi(k)v_{k},h]=[v_{k},h]=v_{k}-\rho(h)v_{k}.$
On the other hand, $[k,h]\in K$ (since $K$ is normal) and $[v_{k},h]\in V$ (since $V$ is normal), so this element belongs to $K\cap V$ and is thus trivial. We see that $v_{k}$ is fixed by all of $H$. We know that $\overline{\pi(K)}$ acts trivially on all of $V$, so $v_{k}$ is fixed by $H$ as well as by $\overline{\pi(K)}$, hence by all of $L$. But by definition of Abels–Noskov groups, the only vector in $V$ fixed by $L$ is the zero vector. This means that $v_{k}=0$, and so $k=\pi(k)\in\ker\rho\leqslant L$. ∎
We are now ready to complete the case $\ker f\cap V$ is trivial. Observe that, since $\rho$ admits no non-zero fixed vectors and $S$ is perfect, a direct computation shows that the commutator subgroup of $(S\times A)\ltimes V$ is $S\ltimes V$.
Lemma 6.8.
Let $G=(S\times A)\ltimes_{\rho}V$ be an Abels–Noskov group, and let $f:G\to H$ be an open epimorphism such that $\ker f\cap V$ is trivial. Let $f^{\mathrm{ab}}:A\to H/f(S\ltimes V)$ be the map induced by $f$. If $h_{1},\dots,h_{n}\in H$ are generators with $n\geqslant\mathrm{d}(G)$, then they admit generating lifts to $G$ via $f$ if and only if $h_{1}f(S\ltimes V),\dots,h_{n}f(S\ltimes V)$ admit generating lifts to $A$ via $f^{\mathrm{ab}}$. If $H$ is Hausdorff, then so is $H/f(S\ltimes V)$.
Proof.
It is clear that, if $h_{1},\dots,h_{n}$ admit generating lifts via $f$, then $h_{1}f(S\ltimes V),\dots,h_{n}f(S\ltimes V)$ admit generating lifts to $A$ via $f^{\mathrm{ab}}$. Thus, suppose $h_{1},\dots,h_{n}\in H$ are generators (with $n\geqslant\mathrm{d}(G)$) such that $h_{1}f(S\ltimes V),\dots,h_{n}f(S\ltimes V)$ admit generating lifts to $A$ via $f^{\mathrm{ab}}$. Set $L=S\times A$ and let $(\ell_{1},v_{1}),\dots,(\ell_{n},v_{n})\in L\ltimes V$ be arbitrary lifts of $h_{1},\dots,h_{n}$. Set $\Delta=\ker f$ and
$F=\overline{\left\langle(\ell_{1},v_{1}),\dots,(\ell_{n},v_{n})\right\rangle},$
so that $F\Delta$ is dense in $G$. What we need to show is that there are $\delta_{1},\dots,\delta_{n}\in\Delta$ such that $(\ell_{1},v_{1})\delta_{1},\dots,(\ell_{n},v_{n})\delta_{n}$ generate $G$.
By the previous lemma, we know $\Delta$ is contained in $L$ and acts trivially on $V$. Set $\bar{L}=L/\overline{\Delta}$, so that $G/\overline{\Delta}$ is naturally isomorphic with $\bar{G}\coloneqq\bar{L}\ltimes_{\bar{\rho}}V$, where $\bar{\rho}:\bar{L}\to\mathrm{GL}(V)$ is the representation induced by $\rho$. Then $\bar{G}$ too is Abels–Noskov. Let ¯$\bar{\ell}_{i}$ be the image of $\ell_{i}$ in $\bar{L}$. By assumption, (¯$(\bar{\ell}_{1},v_{1}),\dots,(\bar{\ell}_{n},v_{n})$ generate $\bar{G}$, so
$\alpha_{\boldsymbol{v}}(\ker\varepsilon_{\boldsymbol{\bar{\ell}}})=V$
by 6.4, where $\alpha_{\boldsymbol{v}},\varepsilon_{\bar{\boldsymbol{\ell}}}$ are defined as usual.
Observe that, since $\overline{\Delta}$ acts trivially on $V$, we have that $\rho(L)=\bar{\rho}(\bar{L})$, so that $\mathbb{R}[\rho]=\mathbb{R}[\bar{\rho}]$ (where this is a genuine equality of subsets of $\mathrm{End}_{\mathbb{R}}(V)$, not just a natural isomorphism). This means that $\varepsilon_{\boldsymbol{\bar{\ell}}}=\varepsilon_{\boldsymbol{\ell}}$ and hence $\ker\varepsilon_{\boldsymbol{\bar{\ell}}}=\ker\varepsilon_{\boldsymbol{\ell}}$, so that
$\alpha_{\boldsymbol{v}}(\ker\varepsilon_{\boldsymbol{\ell}})=\alpha_{\boldsymbol{v}}(\ker\varepsilon_{\boldsymbol{\bar{\ell}}})=V.$
Observe that $f:L\ltimes V\to H$ and $L\longrightarrow\bar{L}$ give rise to the same abelianisation map $f^{\mathrm{ab}}:A\to H/f(S\ltimes V)$. Thus, by 5.2, there are $\delta_{1},\dots,\delta_{n}$ such that $\ell_{1}\delta_{1},\dots,\ell_{n}\delta_{n}$ generate $L$. Since $\Delta$ acts trivially on $V$, we have that $\varepsilon_{\boldsymbol{\ell\delta}}=\varepsilon_{\boldsymbol{\ell}}$. Therefore, by 6.4, the elements $(\ell_{1}\delta_{1},v_{1}),\dots,(\ell_{n}\delta_{n},v_{n})$ generate $G$, as needed.
If $H$ is Hausdorff, then $\ker f\subseteq L$ is closed, so $\ker f^{\mathrm{ab}}$ is closed as well (by 5.2). ∎
6.5 The General Case
Proposition 6.9.
Let $G=(S\times A)\ltimes_{\rho}V$ be an Abels–Noskov group, and let $f:G\to H$ be an open epimorphism. Let $f^{\mathrm{ab}}:A\to H/f(S\ltimes V)$ be the induced epimorphism. If $h_{1},\dots,h_{n}\in H$ are generators with $n\geqslant\mathrm{d}(G)$, then they admit generating lifts to $G$ via $f$ if and only if $h_{1}f(S\ltimes V),\dots,h_{n}f(S\ltimes V)$ admit generating lifts to $A$ via $f^{\mathrm{ab}}$.
If $H$ is Hausdorff, then so is $H/f(S\ltimes V)$.
Proof.
Let $h_{1},\dots,h_{n}\in H$ be generators such that $h_{1}f(S\ltimes V),\dots,h_{n}f(S\ltimes V)$ admit generating lifts to $A$ via $f^{\mathrm{ab}}$. Denote $\Delta=\ker f$. Let $g_{1},\dots,g_{n}\in G$ be arbitrary lifts of $h_{1},\dots,h_{n}$, so that $\left\langle g_{1},\dots,g_{n}\right\rangle\Delta$ is dense in $G$.
The idea of the proof is very simple: we know how to do this for kernels contained in $V$, and we know how to do this for kernels intersecting $V$ trivially; so we first divide by the intersection with $V$, which is a submodule and hence a closed subgroup, and then we divide by the rest, which intersects the image of $V$ trivially. We now spell this out in detail.
Consider $U\coloneqq\Delta\cap V$. It is a normal subgroup of $G$ contained in $V$, so it is a submodule of $V$. In particular, it is closed. Therefore, $G/U$ is naturally isomorphic with $L\ltimes_{\overline{\rho}}W$, $W\subseteq V$ is some submodule complementing $U$, i.e. such that $V=U\oplus W$, and $\bar{\rho}=\rho\!\restriction_{W}$.
Since $\left\langle g_{1},\dots,g_{n}\right\rangle\Delta$ is dense in $G$, we get that $\left\langle g_{1}U,\dots,g_{n}U\right\rangle\hat{\Delta}$ is dense in $G/U$, where $\hat{\Delta}$ is the image of $\hat{\Delta}$ in $G/U$. Clearly, $\hat{\Delta}\cap W$ is trivial. Now, $G/U=L\ltimes_{\overline{\rho}}W$ is still an Abels–Noskov group, and its abelianisation is $A$. Moreover, the map $\bar{f}:G/U\to(G/U)/\hat{\Delta}$ gives rise to the same map $f^{\mathrm{ab}}:A\to H/f(S\ltimes V)$. Thus, since $h_{1}f(S\ltimes V),\dots,h_{n}f(S\ltimes V)$ admit generating lifts to $A$, we may use 6.8 and get that there are $\bar{\varepsilon}_{1},\dots,\bar{\varepsilon}_{n}\in\hat{\Delta}$ such that $g_{1}U\cdot\bar{\varepsilon}_{1},\dots,g_{n}U\cdot\bar{\varepsilon}_{n}$ generate $G/U$. This means that there are $\varepsilon_{1},\dots,\varepsilon_{n}\in\Delta$ such that $g_{1}\varepsilon_{1}U,\dots,g_{n}\varepsilon_{n}U$ generate $G/U$. Therefore, by 6.6 (since $n\geqslant\mathrm{d}(G)\geqslant\mathrm{d}_{\mathbb{R}[\rho]}(V)+1$), there are $u_{1},\dots,u_{n}\in U\subseteq\Delta$ such that $g_{1}\varepsilon_{1}u_{1},\dots,g_{n}\varepsilon_{n}u_{n}$ generate $G$, as needed.
If $H$ is Hausdorff, then $H/f(S\ltimes V)$ is Hausdorff as well by 6.8 (applied to $G/U$, which is Hausdorff in any case). ∎
7 General Connected Lie Groups
Theorem 7.1 ([AN24, Corollary 5.8]).
Let $G$ be a connected Lie group. Then it admits a non-generating subgroup $B\trianglelefteqslant G$, an Abels–Noskov group $L\ltimes_{\rho}V$ and a finite covering map $L\ltimes_{\rho}V\to G/B$.
Lemma 7.2.
Let $G$ be a connected Lie group. Let $B$ be a non-generating subgroup $B\trianglelefteqslant G$ such that there is an Abels–Noskov group $(S\times A)\ltimes_{\rho}V$ and a finite covering map $(S\times A)\ltimes_{\rho}V\longrightarrow G/B$. Then $(G/B)^{\prime}$ is a closed subgroup, and the induced maps
$G/\overline{G^{\prime}}\longrightarrow(G/B)/(G/B)^{\prime}\longleftarrow A$
are finite coverings. In particular, $A$ is isomorphic to $G/\overline{G^{\prime}}$.
Proof.
Since $\rho$ admits no nonzero fixed vectors and $S$ is perfect, a straightforward computations shows that the commutator subgroup of $(S\times A)\ltimes_{\rho}V$ is $S\ltimes_{\rho}V$. In particular, it is closed, so its image in $G/B$ is closed, since the map $(S\times A)\ltimes_{\rho}V\longrightarrow G/B$ is a closed map (being a finite covering). But its image is exactly $(G/B)^{\prime}$ (being an epimorphism), so $(G/B)^{\prime}$ is closed. It is easy to see in general that, if $f:H_{1}\to H_{2}$ is a finite covering, then $(H_{1}/H_{1}^{\prime})\longrightarrow(H_{2}/H_{2}^{\prime})$ is a finite covering, so $A\longrightarrow(G/B)/(G/B)^{\prime}$ is a finite covering as well.
The image of $G^{\prime}$ in $G/B$ is equal to $(G/B)^{\prime}$, which is closed, so it is equal to the image of $\overline{G^{\prime}}$ in $G/B$. The fact the image of $\overline{G^{\prime}}$ is closed in $G/B$ means that $B\overline{G^{\prime}}$ is closed in $G$, which means that the image of $B$ in $G/\overline{G^{\prime}}$ is closed. Now, by [AN24, Lemma 2.4], the image of $B$ in $G/\overline{G^{\prime}}$ is non-generating, so it is a torsion subgroup; being closed, it must be finite. Since it is the kernel of the map $G/\overline{G^{\prime}}\longrightarrow(G/B)/(G/B)^{\prime}$, we get that this map is a finite covering. ∎
Theorem 7.3.
Let $G$ be a connected Lie group, $f:G\to H$ an open epimorphism. Let $f^{\mathrm{ab}}:G/\overline{G^{\prime}}\to H/f\left(\overline{G^{\prime}}\right)$ be the map induced by $f$. If $h_{1},\dots,h_{n}\in H$ are generators with $n\geqslant\mathrm{d}(G)$, then they admit generating lifts to $G$ via $f$ if and only if $h_{1}f\left(\overline{G^{\prime}}\right),\dots,h_{n}f\left(\overline{G^{\prime}}\right)$ admit generating lifts to $G/\overline{G^{\prime}}$ via $f^{\mathrm{ab}}$.
Thus, if $\ker f$ is topologically-finitely-generated (in particular, if $H$ is Hausdorff) and $n\geqslant 2\dim(G/\overline{G^{\prime}})-\dim T$ (where $T$ is the maximal torus of the abelianisation $G/\overline{G^{\prime}}$), then $h_{1},\dots,h_{n}$ admit generating lifts.
Remark 7.4.
In this case, it is not necessarily true that, if $H$ is Hausdorff, then $H/f\left(\overline{G^{\prime}}\right)$ must be Hausdorff as well; see 8.2.
Proof.
Fix a non-generating subgroup $B\trianglelefteqslant G$ such that there is an Abels–Noskov group $(S\times A)\ltimes_{\rho}V$ and a finite covering map $\pi:(S\times A)\ltimes_{\rho}V\to G/B$. Set $\Delta=\ker f$, and consider the following commutative diagram:
[…]
Observe that, even if $H$ is Hausdorff, it is possible that some of the groups in the diagram are not (e.g., $H/f\left(\overline{G^{\prime}}\right)$).
The maps $\alpha,\beta,\gamma$ are the topological abelianisation maps – the quotient by the closure of the commutator subgroup (which happen to be equal to the abstract commutator subgroup in the case of the latter two). The corresponding maps to their right, $\bar{\alpha},\bar{\beta},\bar{\gamma}$, are abelianisation maps of sorts: quotients by subgroups that are somewhere between the abstract commutator subgroup and its closure.
The leftmost vertical maps, $\pi,p,q,\vartheta$, are all absolutely Gaschütz (as are in fact the rightmost vertical maps); the maps $\pi,p,q$ are even finite covering maps. We therefore get that the maps $f,\hat{f},\tilde{f},\tilde{f}^{\mathrm{ab}},\hat{f}^{\mathrm{ab}},f^{\mathrm{ab}}$ are all ‘Gaschütz equivalent’, so $h_{1},\dots,h_{n}\in H$ admit generating lifts via $f$ if and only if $h_{1}f\left(\overline{G^{\prime}}\right),\dots,h_{n}f\left(\overline{G^{\prime}}\right)$ admit generating lifts via $f^{\mathrm{ab}}$.
To spell it out: let $h_{1},\dots,h_{n}\in H$ be generators. Since $\vartheta$ is absolutely Gaschütz, they admit generating lifts to $G$ via $f$ if and only if $h_{1}f(B),\dots,h_{n}f(B)$ admit generating lifts to $G/B$ via $\hat{f}$. Take some arbitrary lifts $x_{1},\dots,x_{n}\in\left((S\times A)\ltimes_{\rho}V\right)/\pi^{-1}(\vartheta(\Delta))$ of $h_{1}f(B),\dots,h_{n}f(B)$; since $\pi:(S\times A)\ltimes_{\rho}V\to G/B$ is absolutely Gaschütz, we get that $x_{1},\dots,x_{n}$ admit generating lifts to $(S\times A)\ltimes_{\rho}V$ via $\tilde{f}$ if and only if $h_{1}f(B),\dots,h_{n}f(B)$ admit generating lifts to $G/B$ via $\hat{f}$. Now, denote the images of $x_{1},\dots,x_{n}$ in $A/\gamma(\pi^{-1}(\vartheta(\Delta)))$ by $\bar{x}_{1},\dots,\bar{x}_{n}$; by 6.9, $x_{1},\dots,x_{n}$ admit generating lifts to $(S\times A)\ltimes_{\rho}V$ via $\tilde{f}$ if and only if $\bar{x}_{1},\dots,\bar{x}_{n}$ admit generating lifts to $A$ via $\tilde{f}^{\mathrm{ab}}$. Denote the images of $\bar{x}_{1},\dots,\bar{x}_{n}$ in $(H/f(B))/\hat{f}((G/B)^{\prime})$ by $y_{1},\dots,y_{n}$. Then, since $p$ is absolutely Gaschütz, $\bar{x}_{1},\dots,\bar{x}_{n}$ admit generating lifts to $A$ via $\tilde{f}^{\mathrm{ab}}$ if and only if $y_{1},\dots,y_{n}$ admit generating lifts to $G/B$ via $\hat{f}^{\mathrm{ab}}$. By the commutativity of the diagram, $y_{1},\dots,y_{n}$ are the images of $h_{1}f\left(\overline{G^{\prime}}\right),\dots,h_{n}f\left(\overline{G^{\prime}}\right)$ under the corresponding map, so, since $q$ is absolutely Gaschütz, $y_{1},\dots,y_{n}$ admit generating lifts via $\hat{f}^{\mathrm{ab}}$ if and only if $h_{1}f\left(\overline{G^{\prime}}\right),\dots,h_{n}f\left(\overline{G^{\prime}}\right)$ admit generating lifts via $f^{\mathrm{ab}}$. ∎
This completes the proof of Theorem 1.3, and of its immediate consequence Theorem 1.2.
8 The Gaschütz Rank
Recall the definition of the Gaschütz rank of an epimorphism and of a group, both appearing in 1.4 in the introduction. In this terminology, 7.3 implies the following:
Corollary 8.1.
Let $G$ be a connected Lie group, $f:G\to H$ an open epimorphism. Let $f^{\mathrm{ab}}:G/\overline{G^{\prime}}\to H/f\left(\overline{G^{\prime}}\right)$ be the map induced by $f$. Then $\text{\lightning}(f)=\max\{\mathrm{d}(G),\text{\lightning}(f^{\mathrm{ab}})\}.$
We always have $\text{\lightning}(f^{\mathrm{ab}})\leqslant 2\dim(G/\overline{G^{\prime}})-\dim T$, where $T$ is the maximal torus of $G/\overline{G^{\prime}}$. If $A$ is a connected abelian Lie group, then $\text{\lightning}(A)=1$ if $A$ is compact (e.g., by [CG18]) and $\text{\lightning}(A)=2\dim A-\dim T$ if $A$ is non-compact, where $T$ is its maximal torus (by 3.9 and 3.10).
If $G$ is any connected Lie group, then
$\mathrm{d}(G)\leqslant\text{\lightning}(G)\leqslant\max\left\{\mathrm{d}(G),2\dim G/\overline{G^{\prime}}-\dim T\right\},$
where $T$ is the maximal torus of $G/\overline{G^{\prime}}$.
Clearly, if $\mathrm{d}(G)\geqslant 2\dim G/\overline{G^{\prime}}-\dim T$ (for example, if $G$ is perfect) then $\text{\lightning}(G)=\mathrm{d}(G)$. If $2\dim G/\overline{G^{\prime}}-\dim T$ is strictly larger than $\mathrm{d}(G)$, then $\text{\lightning}(G)$ may be equal to either end of the spectrum. Its value can be determined exactly by our results in most cases:
  1. 1.
    If $G$ is compact, then we have $\text{\lightning}(G)=\mathrm{d}(G)$ (by [CG18]).
  2. 2.
    If $G/\overline{G^{\prime}}$ is non-compact, then we have $\text{\lightning}(G)=\max\left\{\mathrm{d}(G),2\dim G/\overline{G^{\prime}}-\dim T\right\}$. This is because $\text{\lightning}(G/\overline{G^{\prime}})=2\dim(G/\overline{G^{\prime}})-\dim T$ (by 3.9 and 3.10), and clearly $\text{\lightning}(G)\geqslant\max\{\mathrm{d}(G),\text{\lightning}(G/\overline{G^{\prime}})\}$.
  3. 3.
    If every Hausdorff quotient of $G$ satisfies the condition that its abstract commutator subgroup is closed, then $\text{\lightning}(G)=\max\{\mathrm{d}(G),\text{\lightning}(G/G^{\prime})\}$. [^6] For instance, every Abels–Noskov group satisfies this condition (and hence every group which is Abels–Noskov up to finite coverings), by 6.9. Clearly, $\text{\lightning}(G)\geqslant\max\{\mathrm{d}(G),\text{\lightning}(G/\overline{G^{\prime}})\}$. Conversely, if $f:G\to H$ is an open epimorphism onto a Hausdorff group $H$, we have $\text{\lightning}(f)=\max\{\mathrm{d}(G),\text{\lightning}(f^{\mathrm{ab}})\}$ for the induced map $f^{\mathrm{ab}}:G/G^{\prime}\to H/H^{\prime}$; since $H/H^{\prime}$ is Hausdorff, we have $\text{\lightning}(f^{\mathrm{ab}})\leqslant\text{\lightning}(G/G^{\prime})$.
  4. 4.
    In particular, if every Hausdorff quotient of $G$ has a closed commutator subgroup and $G/G^{\prime}$ is compact, then $\text{\lightning}(G)=\mathrm{d}(G)$ (since $\text{\lightning}(G/G^{\prime})=1$).
We are left with the case $G$ is non-compact, $G/\overline{G^{\prime}}$ is a torus with $\dim(G/\overline{G^{\prime}})>\mathrm{d}(G)$, and $f:G\to H$ is a an open epimorphism onto a group $H$ such that $H^{\prime}$ is not closed. The following example shows that the behaviour of the Gaschütz rank may be a bit tricky in this case.
Example 8.2.
Let $S$ be a Lie group such that $Z(S)$ is non-compact. Let $s\in Z(S)$ be an element of infinite order such that $\left\langle s\right\rangle\cong\mathbb{Z}$ is discrete. Let $T=(\mathbb{R}/\mathbb{Z})^{n}$ (for any $n\geqslant 1$), and let $\vartheta_{i}\in T$ be the element with $\sqrt{2}$ in the $i^{\text{th}}$ coordinate and 00 in the rest (for $i=1,\dots,n$). Let $G=S\times T$, and let $K=\left\langle(s,\vartheta_{1}),\dots,(s,\vartheta_{n})\right\rangle$. Let $f:G\to G/K$ be the quotient map.
Claim.
We have $\text{\lightning}(f)\geqslant n$, and hence $\text{\lightning}(G)\geqslant n$.
Given this claim, we have $\mathrm{d}(G)\leqslant\mathrm{d}(S)+1$, so, if $n$ is large enough, we get that $\text{\lightning}(G)>\mathrm{d}(G)$. If $S$ is semisimple (e.g., $S=\widetilde{\mathrm{SL}_{2}(\mathbb{R})}$, the universal cover of $\mathrm{SL}_{2}(\mathbb{R})$), then $\mathrm{d}(G)=2$, $\text{\lightning}(G/\overline{G^{\prime}})=\text{\lightning}(T)=1$ and $\text{\lightning}(G)=n$. In particular, the formula $\text{\lightning}(G)=\max\left\{\mathrm{d}(G),\text{\lightning}(G/\overline{G^{\prime}})\right\}$ is false in general.
Proof of claim.
We will find $n-1$ generators of $G/K$ that do not admit generating lifts to $G$.
First, observe that $f(S)$ is dense in $G/K$, since $f(S)=f(S\cdot K)$ and $S\cdot K$ is dense in $G$. We may assume $\mathrm{d}(S)\leqslant n-1$, or else the claim is trivial (since $\text{\lightning}(f)\geqslant\text{\lightning}(G)\geqslant\mathrm{d}(G)\geqslant\mathrm{d}(S)$). Let $s_{1},\dots,s_{n-1}$ be generators of $S$, and consider $(s_{1},0),\dots,(s_{n-1},0)\in G$. These descend to generators of $G/K$, since they generate a subgroup that is dense in $S$, and $f(S)$ is dense in $G/K$. Assume by contradiction there are $(\varepsilon_{1}^{S},\varepsilon_{1}^{T}),\dots,(\varepsilon_{n-1}^{S},\varepsilon_{n-1}^{T})\in K$ such that $(s_{1}\varepsilon_{1}^{S},\varepsilon_{1}^{T}),\dots,(s_{n-1}\varepsilon_{n-1}^{S},\varepsilon_{n-1}^{T})$ generate $G$. Then, in particular, $\varepsilon_{1}^{T},\dots,\varepsilon_{n-1}^{T}\in\left\langle\vartheta_{1},\dots,\vartheta_{n}\right\rangle$ generate a dense subgroup of $T$, which is impossible.
If $S$ is semisimple, then $\mathrm{d}(S)=2$ and $\mathrm{d}(G)=2$. The fact $\text{\lightning}(G)\leqslant n$ follows from 7.3, so $\text{\lightning}(G)=n$, since we just proved $n\leqslant\text{\lightning}(G)$. ∎
Of course, if $S/\overline{S^{\prime}}$ is non-compact, we are simply in Case 2, where $G/\overline{G^{\prime}}$ is non-compact and hence the formula $\text{\lightning}(G)=\max\left\{\mathrm{d}(G),2\dim G/\overline{G^{\prime}}-\dim T\right\}$ holds (where in this case $T=G/\overline{G^{\prime}}$). However, if $S/\overline{S^{\prime}}$ is compact, we are in a scenario that is not covered by any of the cases above. The simplest example of such an $S$ is $\widetilde{\mathrm{SL}_{2}(\mathbb{R})}$, but it is also possible for such an $S$ to be an algebraic group:
Example 8.3.
Let $\omega:\mathbb{R}^{2}\times\mathbb{R}^{2}\to\mathbb{R}$ denote the determinant form:
$\omega(x,y;x^{\prime},y^{\prime})=x\cdot y^{\prime}-x^{\prime}\cdot y.$
Let $S=\mathrm{SL}_{2}(\mathbb{R})$, which acts on $\mathbb{R}^{2}$ by matrix multiplication, and preserves $\omega$ (i.e., $\omega(sv,su)=\omega(u,v)$ for every $s\in S$, $u,v\in\mathbb{R}^{2}$). In particular, the $S$-module $\mathbb{R}^{2}$ is irreducible. Now, consider the Heisenberg group $H$, defined to be the set of triples $(x,y,z)\in\mathbb{R}^{3}$ with the group law given by:
$(x,y,z)\cdot(x^{\prime},y^{\prime},z^{\prime})=\left(x+x^{\prime},y+y^{\prime},z+z^{\prime}+\frac{1}{2}\omega(x,y;x^{\prime},y^{\prime})\right).$
This is a connected Lie group. The standard representation of $S$ on $\mathbb{R}^{2}$ induces an action on $H$ defined by:
$A.(x,y,z)=\left(A.(x,y),z\right).$
That is, $S$ acts only on the first two coordinates. The fact that this is indeed a group action follows from the fact $S$ preserves $\omega$.
Now, we form the semi-direct product $G=S\ltimes H$, often referred to as the ($1$-dimensional) Schrödinger group [^7] This is the group of symmetries of the Schrödinger equation of a free particle in a $1$-dimensional space.. It is clear that $S$ is a real algebraic group. We claim that the centre of $G$ is non-compact, and that $G$ is perfect.
Proof.
It is easy to see that the subgroup $Z$ consisting of all triples $(x,y,z)$ with $x=y=0$ is in the centre of $H$, and that this subgroup is isomorphic to $\mathbb{R}$. Since $G$ acts on this subgroup trivially, $Z$ is central in $G$ as well. Therefore, the centre of $G$ is non-compact.
Now, it is not hard to see from the group law that the commutator subgroup of $H$ is contained in $Z\cong\mathbb{R}$. It is moreover connected and non-trivial, and so it must in fact be equal to $Z$. In particular $Z$ is contained in the commutator subgroup of $G$. It thus suffices to show that the quotient of $G$ by $Z$ is perfect. Now, $G/Z$ is isomorphic to the group $S\ltimes\mathbb{R}^{2}$ with the standard action. The commutator $[S,\mathbb{R}^{2}]$ is a subgroup of $\mathbb{R}^{2}$, thus an $S$-submodule, because $S$ is connected. But the action of $S$ on $\mathbb{R}^{2}$ is irreducible, and so $[G,\mathbb{R}^{2}]=\mathbb{R}^{2}$. Since $S$ is perfect, we see that $S\ltimes\mathbb{R}^{2}$ is perfect as well. ∎
Remark 8.4.
While writing this paper, it was useful to have in mind another invariant, the Gaschütz-* rank of $G$, $\text{\lightning}^{*}(G)$. It is defined to be the minimal $m\in\mathbb{N}\cup\left\{\infty\right\}$ such that every open epimorphism $f:G\to H$ whose kernel is topologically-finitely-generated satisfies $\text{\lightning}(f)\leqslant m$. This invariant is better behaved with quotients: if $N$ is a normal non-generating subgroup of $G$ which is topologically-finitely-generated (and recall that, in Lie groups, every closed normal subgroup is topologically-finitely-generated), we have $\text{\lightning}^{*}(G)=\text{\lightning}^{*}(G/N)$. In particular, for every connected Lie group one has
$\text{\lightning}^{*}(G)=\max\{\mathrm{d}(G),2\dim(G/\overline{G^{\prime}})-\dim T\},$
where $T$ is the maximal torus of $G/\overline{G^{\prime}}$.